### v2

# Evaluating Generative Ad Hoc Information Retrieval

Abstract.
Recent advances in large language models have enabled the development of viable generative retrieval systems. Instead of a traditional document ranking, many generative retrieval systems directly return a grounded generated text as an answer to an information need expressed as a query or question. Quantifying the utility of the textual responses is essential for appropriately evaluating such generative ad hoc retrieval. Yet, the established evaluation methodology for ranking-based retrieval is not suited for reliable, repeatable, and reproducible evaluation of generated answers. In this paper, we survey the relevant literature from the fields of information retrieval and natural language processing, we identify search tasks and system architectures in generative retrieval, we develop a corresponding user model, and we study its operationalization. Our analysis provides a foundation and new insights for the evaluation of generative retrieval systems, focusing on ad hoc retrieval.
1. Introduction
The development of large language models (LLMs) has prompted search engine and AI companies to innovate the way search results are presented: using LLMs to directly generate a textual answer that satisfies an information need. While LLMs can generate unreliable information (Alkaissi and McFarlane, 2023; Ji et al., 2023; Koopman and Zuccon, 2023), conditioning their inference on relevant documents has emerged as a potential technique to ground their generated statements (Lewis et al., 2020; Mialon et al., 2023). This may relieve users of the (cognitive) effort of acquiring the needed information from individual search results themselves, which affords a change in the design of a search engine results page (SERP; Figure 1): instead of the proverbial list of “ten blue links” (list SERP, left), a generated text with references is shown (text SERP, right). The first public prototypes of this kind were You.com’s You Chat and Neeva AI, closely followed by Microsoft’s Bing Copilot, Google’s Gemini, Perplexity.ai, Baidu’s Ernie,111See https://chat.you.com; Neeva has shutdown; https://chat.bing.com; https://gemini.google.com; https://perplexity.ai; https://yiyan.baidu.com. and other research prototypes (Koopman et al., 2023; Zhang and Pradeep, 2023).
On the left, a schematic illustration of a list of search results as it is shown on conventional search results pages. On the right, a schematic text with citations and a reference list of corresponding links below it as a novel presentation of a search result.
Far ahead of this development, Sakai et al. (2011) raised an important question: How can search engines that use text SERPs be evaluated? Evaluating text SERPs is not straightforward, since the modern theory and practice of evaluation in information retrieval is premised on the assumption that search results are presented as list SERPs.222Extensive research on search interfaces has included many alternatives, search features, interaction designs, and result visualizations (Hearst, 2009; Wilson, 2011; Liu et al., 2021). Nonetheless, with the growth of the Web, the list SERP design became a de facto standard for web search.
According to list SERP user models, a ranked list of results triggers a certain user behavior like reading the results in order until the information need is satisfied or the search is abandoned. In decades of research, a comprehensive theoretical framework of reliable and validated evaluation methods has been built to assess the quality of document rankings with respect to information needs. Replacing ranked results by a generated text undermines this foundation.
In this paper, we focus on questions related to transferring established list SERP evaluation methodology to text SERPs. Our approach is theory-driven and based on a systematic analysis of relevant literature from information retrieval (IR) and related fields. Our contributions relate to the system, user, and evaluation perspectives. Starting with a definition of what generative ad hoc retrieval is, we distinguish two fundamental system models for generative retrieval and contextualize them in Broder’s ((2002)) taxonomy of search tasks (Section 2). We then devise a user model for text SERPs, grounded in related behavioral studies (Section 3). Finally, we revisit IR evaluation methodologies to develop a foundation for text SERP effectiveness measures and for the reliable evaluation of generative ad hoc retrieval (Section 4).
2. Generative Retrieval Tasks
In this section, we define the task of generative ad hoc retrieval, we review the two fundamental paradigms of its operationalization, we discuss its contribution on top of traditional ad hoc retrieval, and we distinguish it from other generative retrieval tasks.
2.1. Generative Ad Hoc Retrieval
The tasks of language generation and of retrieval appear to be distinct. However, retrieval systems and generative language models are both created using large collections of documents (see Figure 2 top), and the usefulness of both depends on tuning them with user needs, expressed as queries or prompts . The users of an IR system want to retrieve the most relevant documents for a query, and the users of a generative language model want to generate the most helpful text for a prompt. From an IR perspective, the most salient difference between retrieval and generation is as follows: a retrieval model induces a ranking on a finite document collection , while a generative language model induces a ranking on the infinite set of all possible texts . In practice, retrieval models are used to return the top- ranked documents, and generative language models to return just one of the many possible relevant texts from . Generative models have therefore recently been framed as infinite indexes (Deckers et al., 2023).
As a retrieval model can only return existing documents, the information available in determines the degree to which a user’s information need can be satisfied. Still, the user has to examine the returned documents for the desired information. A generative language model instead attempts to alleviate the effort of examining documents by returning a tailored response that compiles all desired information. Yet, the factual accuracy of current generative language models is often lacking and prone to confabulations or hallucinations (Zhao et al., 2023; Koopman and Zuccon, 2023; Ji et al., 2023; Alkaissi and McFarlane, 2023) (i.e., there is only a very small subset of accurate texts among all possible texts ).333For cases like counterfactual information needs (e.g., “What if Columbus didn’t settle America?” (Kiciman et al., 2023)), strong confabulation capabilities could be explicitly desirable, though.
Generative ad hoc retrieval can be described as combining retrieval and generation in ad hoc scenarios (one query, one result) in a way that the used systems’ respective advantages complement each other. The two fundamental combination approaches are (Figure 2, bottom): retrieving relevant documents from and generating an answer from them, or generating a response and “verifying” its statements by retrieving supporting documents from .
2.2. Operationalization Paradigms
System for generative ad hoc retrieval require two components (Gao et al., 2023): (1) retrieval, to gather existing documents from a collection for a query, and (2) generation, to generate a text for a prompt. Both components can be combined in two ways (Gao et al., 2023): retrieval-then-generation and generation-then-retrieval, which form the two paradigms of operationalizing generative ad hoc retrieval (Figure 2, bottom). In a retrieval-then-generation approach, a language model is conditioned with retrieved source material, for instance, by adding evidence to its input prompt (Izacard and Grave, 2021; Khot et al., 2023; Lazaridou et al., 2022; Shi et al., 2023), attending to retrieved sources during inference (Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022), chaining ideas (Jiang et al., 2022), or iterative self-attention starting from sources (Zhang et al., 2021b). In a generation-then-retrieval approach, the retrieval model is used to find sources for generated text passages. Though this approach has received little attention (AutoGPT Contributors, 2023), it resembles retrieving references for individual generated statements, similar to claim verification (Wadden et al., 2020).
With increasing inference speeds of generative language models, arbitrarily ordered combinations of the retrieval and the generation step are possible, leading to multi-step generative ad hoc retrieval. The simplest form might be iterative cyclical patterns like generating a text passage that is used as a query to retrieve relevant sources, which in turn serve as context for the next iteration, etc. Applications are the continuous generation of text (Ram et al., 2023; Jiang et al., 2023; Semnani et al., 2023), retrieving sources in multiple steps (Patwardhan and Yang, 2023), or the refinement of a text through iterative inference (Khattab et al., 2022; AutoGPT Contributors, 2023).
In this paper, we focus on the evaluation of the text SERP output of (multi-step) generative ad hoc retrieval, but do not consider evaluating any step individually.
2.3. Generative (Ad Hoc) Search Tasks
Ad hoc retrieval has a long history with a large body of respective research (Manning et al., 2008). Its goal is to output a single result (ranking) for a single-query input (i.e., the information need must be satisfied without knowing any previous queries or interactions). In 2002, Broder suggested a now well-known taxonomy of search tasks (Broder, 2002) and related them to three generations of web search systems (see Table 1). Each generation utilizes a new source of information in addition to those of its predecessors to meet new user intents. The first generation supports informational tasks, relying only on the information found within single documents to support a user’s intent to acquire (parts of) that information. The second generation additionally exploits document relations, supporting users to reach a specific site, document, or the most authoritative one among many alternatives (i.e., navigational tasks). The third generation blends results from different vertical systems and multimodal results into a single SERP to support a user in performing transactional tasks.
Generative retrieval systems can be seen as a new, 4th generation of web search systems. They enable the synthesis of new “documents” relevant to a user’s information need. Given a sufficiently complex need that cannot be satisfied by returning the URL of a web page, generative retrieval systems are primarily used to synthesize a comprehensive overview of the information required, condensed into a long-form answer. Prior search system generations delegated the condensation of information from multiple sources to the users who then often needed to study several documents from a list SERP to satisfy their needs. Generative models promise to reduce this extra work and cognitive load, so that users need only read one generated text.444Already in (2011), Sakai et al. (2011) have previously proposed to automatically identify relevant information nuggets in retrieved documents and to only present the nuggets in a list, but did not consider the aspect of condensing them. Additionally, the “synthesizing” nature of generative retrieval systems can conceivably be harnessed to generate new pieces of information not contained in the retrieved sources, rendering the generative model itself a source of information.
| Gen. | Search task | Information source | User intent | Year |
| 1st | informational | Document | Acquire | 1995 |
| 2nd | + navigational | + Document relations | + Reach | 1998 |
| 3rd | + transactional | + Search verticals | + Perform | 2002 |
| 4th | + synthetic | + Generative models | + Condense | 2023 |
While the search tasks addressed by generative retrieval systems are in many cases informational in nature, we argue to consolidate them as synthetic search tasks. Typical examples are opinion needs (“Should society invest in renewable energy?”) or decision-making needs (“Should I get life insurance?”). Such needs are not fully supported by the first three system generations: in contrast to (1) informational tasks, the information is likely spread across multiple documents; in contrast to (2) navigational tasks, no single page is premeditated to be reached by the user; and in contrast to (3) transactional tasks, the goal of condensing the information should be addressed on the system side. As if foreseeing synthetic retrieval, Broder explicitly constrained informational queries and first-generation systems to static content: “The purpose of such [informational] queries is to find information assumed to be available on the Web in a static form. No further interaction is predicted, except reading. By static form we mean that the target document is not created in response to the user query.” (Broder, 2002, page 5)
The fourth generation of web search engines supports synthetic search tasks and enables users to access a single, comprehensive generated answer (document) that satisfies a complex need with an in-depth analysis covering multiple perspectives. Although the Web may offer a static (set of) document(s) to satisfy such a need, an ideal generative retrieval system retrieves them, synthesizes missing information, presents a coherent answer, and grounds it in the retrieved sources, i.e., dynamically addresses the query.
2.4. A Taxonomy of Generative Retrieval
A taxonomy graph that has generative information retrieval at the top level, and then is subdivided into generation-augmented retrieval (GAR) and retrieval-augmented generation (RAG) on the second level. The GAR node further consists of indexing-time GAR and query-time GAR, while the RAG node further consists of attention-level RAG and prompt-level RAG.
‘Generative retrieval’ or ’generative IR’ are umbrella terms for a diversity of approaches that use generative models to solve retrieval tasks.555See also the recent SIGIR workshop on generative IR (Bénédict et al., 2023). Following Arora et al. (2023), Figure 3 categorizes these approaches into generation-augmented retrieval (GAR) and retrieval-augmented generation (RAG). Notably, GAR approaches create traditional list SERPs, while RAG approaches generate text SERPs.
In GAR approaches, generative models are used to enhance the traditional search architecture, which can be done at indexing time and at query time. At indexing time, generative models can be used for augmenting documents (Nogueira et al., 2019; Gospodinov et al., 2023; MacAvaney et al., 2020; Formal et al., 2021; Zhuang and Zuccon, 2021) with confabulated or hallucinated content, or for replacing the standard indexing process with what are commonly termed ‘differentiable indices’ by, for instance, generating document identifiers like page titles (De Cao et al., 2021; Thorne, 2022; Chen et al., 2022), URLs (Ziems et al., 2023), or (structured) string identifiers (Zhou et al., 2023; Tay et al., 2022; Zhuang et al., 2022; Wang et al., 2022). At query time, generative models can be used for augmenting queries (MacAvaney et al., 2021; Alaofi et al., 2023), or for modeling relevance by, for instance, generating parts of existing documents from the query and retrieval by string matching (Bevilacqua et al., 2022), predicting a (re-)ranking directly (Sun et al., 2023), or using special tokens as relevance signal (Nogueira et al., 2020; Ma et al., 2023; Qin et al., 2023; Zhuang et al., 2023).
In RAG approaches—the focus of our paper—, generative models are augmented with retrieval capabilities; either internally as what we term ‘attention-level RAG’, where the context attended to during generation is retrieved concurrently (Guu et al., 2020; Jiang et al., 2022; Borgeaud et al., 2022), or externally as ‘prompt-level RAG’, where the retrieved context is inserted into the prompt. Orthogonally, one can distinguish the RAG variants retrieval-then-generation and generation-then-retrieval.
Beyond GAR and RAG, generative models can also be used to directly generate a response without relying on retrieved information (Robinson and Wingate, 2023), i.e., as infinite indexes (Deckers et al., 2023). This may involve generating multiple candidates and selecting the best one or regenerating a new response conditioned on the previous ones (Yu et al., 2023). Moreover, an answer to a generative ad hoc request can also be the first turn of a conversational search (Salton, 1969; Radlinski and Craswell, 2017; Culpepper et al., 2018), where generative models have led to new tools (Zhang et al., 2021a; Miller et al., 2017) and dialog options (Zamani et al., 2020).
3. A User Model for Generative IR
Any IR system should align with user expectations, thus, evaluation needs to be grounded in a user model. Yet, existing user models that have been derived to facilitate evaluation in IR are based on the assumptions of list SERPs. After preliminary considerations to derive a text SERP user model (Section 3.1), we first consider the general search process of a user and explore how it relates to generative approaches (Section 3.2). Then, we follow the evaluation methodology proposed by Agosti et al. (2014): first, we define evaluation objectives (Section 3.3) and then we devise a user model that corresponds to these objectives (Section 3.4). This makes it possible to later derive metrics that operationalize the user model.
The general structure is reflected in Figure 4. We base our proposed evaluation methodology on the ad hoc information search process (Vakkari, 2016) as seen from the users’ perspective (top of the figure), and formulate evaluation objectives that correspond to each component (bottom of the figure), which take into account the evaluation setting from which a user model can be induced. Traditional IR can assist the user only during Selection with a list SERP. Meanwhile, generative retrieval encompasses all three steps of Selection, Interaction, and Synthesis, to support the synthetic search task and respond with a text SERP. An evaluation of a generative retrieval system should therefore focus on these steps, which are mirrored in the Retrieval, Grounding, and Presentation evaluation objectives.
The figure consist of two aligned rows. The upper row illustrates the user information search process, with the four components of Formulation, Selection, Interaction, and Synthesis. The lower row illustrates generative IR evaluation objectives, with the four components Prompting, Retrieval, Grounding, and Presentation. Three pairs of components are related: Selection and Retrieval, Interaction and Grounding, as well as Synthesis and Presentation. These three pairs represent the steps supported by generative IR systems.
3.1. Preliminary Considerations
Evaluation Setting
In traditional retrieval, the user is presented with a ranked list of documents (list SERP), each typically referenced by a linked title, snippet, and URL. In generative IR, instead, a response text is presented (text SERP), i.e., a sequence of statements, each optionally referencing one or more sources of evidence in support of the statement. A statement can be any consecutive passage of text, ranging from a single word or phrase to a sentence and even one or more paragraphs. In this context, statements are considered ‘atomic’ in the sense that we disregard the nesting of statements of different lengths, and that they support one or more claims that are pertinent to the user’s information need. They are comparable to the concept of ‘atomic/semantic content units’ (Liu et al., 2023a; Nenkova et al., 2007) in summarization evaluation, or ‘information nuggets’/‘retrieval unit’ in traditional IR (Dang and Lin, 2007; Sakai et al., 2011; Sakai, 2023; Chen et al., 2023). A statement can be referencing none, one, or more than one source. References explicitly link to a source, like a web document containing the information on which the generated statement is based and by which it is grounded. The evaluation commences ad hoc, i.e., with a single-query and without session-based or conversational elements.
Evaluation Paradigms
To estimate the effectiveness of retrieval systems, offline evaluation within a Cranfield-style evaluation setting (Cleverdon, 1967) is the de facto standard in IR research. It is used to estimate the satisfaction of users with the output of a system for a given topic (query) by relying on a pool of documents judged by assessors for a given topic set (Sanderson, 2010). This pool of relevance judgments can be reused in subsequent experiments by matching retrieved document identifiers with the ones in the pool. Once a pool of relevance judgments has been compile for a given topic, large-scale evaluations of alternative search systems can be performed. However, since the output documents of a generative retrieval system are novel at query time, they cannot be straightforwardly matched with existing pools of relevance judgments. This bears similarities to the unjudged document problem, which is traditionally solved by assuming non-relevance (Fröbe et al., 2023)—a heuristic, which is not feasible for generative IR. Therefore, more sophisticated transfer methods are required to adapt offline evaluation for generative retrieval.
Alternatively, evaluation of generative retrieval systems can be conducted in an online fashion (Sallam et al., 2023). Here, for each run of a system configuration, all its output documents are judged anew, without relying on previous judgments, by collecting user feedback about a system (Kelly, 2009), e.g., by rating their satisfaction. This maximizes the manual effort required to judge runs during experimentation, often happens uncontrolled, is expensive and time-consuming to conduct, and is challenging to replicate, repeat, and reproduce (Renaud and Azzopardi, 2012). To mitigate these problems especially in an academic setting, where access to human user data is often limited, much research went into simulating agents to analyze (interactive) information systems (Maxwell et al., 2015; Maxwell and Azzopardi, 2016; Câmara et al., 2022). However, these agents cannot yet compete with “real” human feedback. Automatic evaluations, where the output of one model is judged by another, has been proposed as a possible way forward (Liu et al., 2023b; Yue et al., 2023), but judging the output of generative models by means of other models has itself been criticized (Sakai, 2023; Bauer et al., 2023; Faggioli et al., 2023).
3.2. Components of the User Search Process
To derive suitable evaluation objectives, first, we have to consider the search process a user undergoes when performing an ad hoc search. Specifically, the synthetic search task enabled by generative retrieval systems should be reflected here. Based on Vakkari (2016), a users’ process encompasses four steps: search formulation, source selection, source interaction, and synthesis of information. Each of these can be mapped to capabilities of generative retrieval systems.
First, during Formulation, the user crafts a specific query that expresses the desired search outcome, addressing their information need. This is no different in generative retrieval systems, though what is called a ‘query’ in IR is a ‘prompt’ in artificial intelligence research. To avoid confusion, we stick to the term ‘query’. In this paper, we leave this step entirely to the user who (iteratively) adapts their search formulation. Yet, we do acknowledge that this task may also be framed as a system task with the goal of enhancing the users’ original query with more context or prompt templates, akin to query suggestion and query expansion in traditional retrieval.
Second, during Selection, the user is presented with a result list and can then examine each entry, possibly through surrogates like snippets. They can assess whether the presented results and their information need align and thus build a focused selection of sources. In generative retrieval systems, this stage corresponds to the system selecting sources that contain potentially relevant information.
Third, during Interaction, the user analyzes the content of each previously selected result in-depth. The aim is to extract and structure the relevant information from each source that addresses the knowledge gap that induces their information need. In generative retrieval systems, this step is supported by the model attending to relevant pieces of information previously retrieved.
Finally, during Synthesis, the user assembles the search outcome. They combine relevant information identified in multiple sources into a coherent answer to their query. In generative IR, this corresponds to the inference of the response text addressing all aspects of the user’s query with information from the previously selected sources. This is key in enabling the synthetic search task. Note that interaction and synthesis often commence concurrently.
3.3. Evaluation Objectives
For each of the components of the search process, we define a corresponding evaluation objective in the context of generative IR. These are not considered evaluation steps, but rather objectives of the evaluation of the system as a whole.
Prompting Objective
Formulation is reflected in the evaluation of the models’ input prompt. While search formulation is an important component to evaluate, we believe it is out of the scope of this paper, since, as previously argued, the formulation step is left to the user. For further reading, the issue of prompt engineering as an emergent field of research is covered in extensive related work (Shin et al., 2020; Reynolds and McDonell, 2021; Gao et al., 2021; Liu and Chilton, 2022; Sorensen et al., 2022; White et al., 2023; Yang et al., 2023).
Retrieval Objective
Selection is reflected in the assessment of the context from which a generative retrieval system draws its information. The retrieved sources (and any relevant information that was not retrieved) directly impact the quality of the generated response. Therefore, the retrieval objective assesses a system’s ability to identify source documents satisfying a user’s information need. This includes its ability to select (1) relevant (aligning with the users’ information need), (2) diverse (covering a variety of information), (3) informative (containing valuable information), and (4) correct (providing accurate information) documents from a collection.
Grounding Objective
Mimicking Interaction, generative ad hoc IR models draw upon reference documents as evidence to generate a response. Yet, grounded text generation may suffer from hallucinations of broadly two types (Maynez et al., 2020): intrinsic hallucinations, where the model wrongly modifies information from the source documents, and extrinsic hallucinations, where the model generates information that is not present in the source documents. Both negatively impact the quality of the generated response (Maynez et al., 2020; Lux et al., 2020). Therefore, the grounding objective assesses a system’s ability to correlate its generated output with information from source documents. This includes its ability to (1) identify (find relevant information), (2) paraphrase (restate that information correctly), and (3) establish consistency (not produce contradictions to other sources).
Presentation Objective
The relevant information across multiple documents has to be synthesized into a single search outcome. This resembles multi-document summarization. Therefore, the presentation objective assesses a systems’ ability to convey information to a user through the generated response in a useful manner, i.e., its ability to produce text that is (1) concise (at a level of granularity sensible given the topic or needed by the user (Dang, 2005)), (2) coherent (in a uniform style), and (3) accessible (written in an understandable way, which, again, is dependent on user needs).
3.4. Components of the User Model
Generative IR poses a challenge for developing a user model. As it is a new IR paradigm, little to no user feedback, A/B tests, laboratory studies, or user behaviour data is available in the academic context from which insights about user behavior may be derived. Further, the information search process that user behavior is traditionally grounded in is replaced (in part) by the generative system. Additionally, the assumptions of traditional user models are made for list SERPs and thus have to be revisited, taking into account the previously established evaluation objectives. To this end, we contribute a user model for generative IR, extrapolating from established evaluation practices in related fields, like question answering, summarization, and traditional IR. We follow the considerations of Carterette (2011), who argues that an IR-focused user model is constituted by three distinct (sub-)models: (1) a utility model (how each result provides utility to the user), which induces a gain function; (2) a browsing model (how the user interacts with results), which induces a discount function; and (3) an accumulation model (how the individual utility of documents is aggregated), combining the individual gain and discount values.
3.4.1. Utility Model for Generative IR
We first motivate a utility model by surveying literature on evaluation in IR and related fields. We identify 10 dimensions of utility applicable to the ad hoc synthetic search task. These are grouped into five top-level categories of Coherence, Coverage, Consistency, Correctness and Clarity. We further distinguish the unit from which gain is derived, as either an individual statement in the response (statement-level) or the response as a whole (response-level). Figure 5 summarizes the proposed dimensions of utility as taxonomy, divided into response- and statement-level dimensions with corresponding objectives marked.
A taxonomy tree growing from left to right which illustrates the utility dimensions in generative ad hoc retrieval. The root node is utility. It subdivides into two sub-levels, where each child on the first has two further childs on the third. Each child on the first level corresponds to an evaluation objective. From top to bottom, the nodes read: 1. coherence, subdivided into stylistic and logical (presentation objective); 2. coverage, subdivided into broad and deep coverage (retrieval objective); 3. consistency, subdivided into internal and external (grounding objective); 4. correctness, subdividing into factual and topical (retrieval objective); and 5. clarity, subdividing into language and content (presentation). Brackets to the right of tree indicate the level each dimension is measured at. Coherence, coverage, and consistency are measured on the response level. External consistency, correctness, and clarity are measured on the statement level.
Coherence
Coherence is a response-level dimension of utility and refers to the manner in which the response is structured and presented. This includes arranging statements to form a coherent narrative without contradictions (Radev and McKeown, 1998; Shah et al., 2021) (Logical Coherence), but also a uniform style of speech (Stylistic Coherence), rendering it readable and engaging (Jin et al., 2020; Capra and Arguello, 2023). Both implement the presentation objective at response level, amounting to “Is the response structured well?” (Logical Coherence) and “Does the response have a uniform style of speech?” (Stylistic Coherence).
Coverage
Coverage measures the cumulative extent to which presented information is pertinent to the users’ information need. It can be subdivided into two forms (Cambazoglu et al., 2021): Broad Coverage, i.e., whether the response covers a breadth of diverse information (Zheng et al., 2012), and Deep Coverage, i.e., whether the response provides in-depth detailed information with high informativeness (Maxwell et al., 2017). Coverage implements the retrieval objective at response level, amounting to “Does the response cover diverse information?”(Broad Coverage) and “Does the response offer detailed information?” (Deep Coverage).
Consistency
A commonly observed problem with source-based text generation is inconsistency (Huang et al., 2021) between source and generated text, which is detrimental to utility. Inconsistencies may also occur across multiple statements within a response, rendering it both a statement-level and response-level dimension. We refer to the first as Internal Consistency (response level), which involves assessing the consistency between statements that constitute the response, ensuring that they do not contradict each other (Nishino et al., 2019; Sakai, 2023; Capra and Arguello, 2023). It should be noted that this does not mean that different conflicting perspectives on a topic can not be reflected in the response, however, these should be explained. The second, External Consistency (statement level), involves assessing the consistency between a statement and its source document(s), ensuring that the generated text aligns in terms of content and context (Maynez et al., 2020; Yue et al., 2023; Sakai, 2023). External inconsistencies are often introduced through model hallucinations (Ji et al., 2023). Consistency is different from factual correctness, as it only assesses the alignment of a statement with the source, and not its objective truth. Both notions implement the grounding objective but on different levels, amounting to “Is the response free of contradictions?” (Internal Consistency) and “Is the statement conveying from sources accurately?” (External Consistency)
Correctness
Correctness gauges to which degree the information provided in the response is factual, reliable, and addressing the user’s information needs. We subdivide correctness into Factual and Topical Correctness. The former captures the degree to which a statement reproduces information that can be assumed as objectively true. Yet, outside of small-scale domain-specific evaluation studies (Sallam et al., 2023) fact-checking remains a challenge (Nakov et al., 2021) and is thus often reduced to a simpler approach, framing it in terms of verifiability (Liu et al., 2023b), not truth. Here, the main requirement is that a piece of information can be attributed to a reliable reference, bestowing it correctness (Wikimedia Foundation, 2023; Yue et al., 2023). Topical correctness denotes whether a statement aligns with the users’ information need (Maddalena et al., 2017; Yang, 2017; Roitero et al., 2018). Both operationalize the retrieval objective at the statement level, amounting to “Does the statement state things that are verifiably true?” (Factual Correctness) and “Does the statement state things within the scope of the user’s information need?” (Topical Correctness).
Clarity
The response given by a generative retrieval system should be expressed in a clear and understandable manner (Zhu et al., 2009; Sameki et al., 2016). This includes the use of language in a concise (Dang, 2005; Sakai, 2023), comprehensible (Cambazoglu et al., 2021), lexically and grammatically correct, and accessible way to the user (Language Clarity). Note that language clarity does not reflect fluency, which is assumed already at human-level for model-generated text (Sakai, 2023), but rather the response being in the appropriate language register. For example, a technical query might warrant an academic style of writing in the response, while a joke question might afford a more jovial tone. Orthogonal to this, the way a statement is written should always clearly communicate the most salient information (Schuff et al., 2022), and where it stems from (Nourani et al., 2019), in order to make the response explainable (Content Clarity). Both operationalize the presentation objective at the statement level, amounting to “Is a statement written in an easily readable way?” (Language Clarity) and “Does the statement put its focus on the most salient points?” (Content Clarity).
3.4.2. Reading Model for Generative IR
For list SERPs, user interaction is modeled by a browsing model, of which two fundamental kinds exist. The set-based kind assumes that a user indiscriminately examines all retrieved documents, while the ranking-based model assumes a user traversing retrieved documents ascending in rank, stopping when either their information need is fulfilled or the search is aborted (Carterette, 2011). Aborting the search is primarily motivated by the effort being too high to justify continuing to browse. Yet, in generative IR, the selection and interaction steps of the search process are supported by the system, thus the user only has to read the generated text. This reduces the effect of stopping criteria grounded in effort, with most users only aborting their search once their knowledge gap is fulfilled, the response is deemed insufficient, or the whole response was read. This is neither set-based, as reading the response is a sequential process and early stopping might occur, nor traditionally ranking-based, as aborting the search is not motivated by effort but rather by search (dis-)satisfaction.
We therefore propose a reading model for generative IR, as an evolution of the standard browsing model, which instead models the attention a user places on each statement while reading. Since there are no empirical studies on reading behavior for generative retrieval at present, we instead turn to related work in reading behavior for document comprehension. We identify a total of six criteria which influence the reading process of documents for an information-seeking purpose, three of which we deem relevant to the case of generative IR. First, Progression (Buscher et al., 2012; Li et al., 2018, 2019; Zheng et al., 2019; Wu et al., 2023) implies that users parse a document sequentially, i.e., progress through the statements constituting the text in order. Second, Decay (Frey et al., 2013; Li et al., 2018, 2019; Zheng et al., 2019; Wu et al., 2023) implies that the reading attention diminishes over the span of the text. Third, Saturation (Li et al., 2018, 2019) implies that users abort once they read enough to fulfill their information need. In sum, this characterizes the browsing behavior of a user as sequentially reading with decaying attention, stopping early if saturated.
While three other characteristics of reading behavior have additionally been found in related work, we deem them superfluous for this reading model: (1) Although perceived relevance is heightened following a relevant statement (Li et al., 2018; Zheng et al., 2019), we adopt the same restriction as a static browsing model for ad hoc retrieval evaluation that neglects inter-statement effects (Moffat et al., 2013, 2015). While the effect is acknowledged, its effect size may not justify the costs of its operationalization. (2) Although attention is highest around query terms (Li et al., 2018; Zheng et al., 2019), we model utility not per token, but at the statement level, thus rendering this effect constant. (3) Although users skip content non-relevant to them (Li et al., 2018; Gwizdka, 2014; Buscher et al., 2012), such statements already receive zero utility.
The properties of the proposed reading model can be related to the (Moffat et al., 2017) framework of browsing models for list SERPs. The conditional ‘continuation probability’ () denotes how likely a user is to continue to browse to the next statement after having seen one. This can alternatively be framed in terms of the ‘weight’ (), which refers to the probability of a user reaching each step of the sequence. The ‘last probability’ (), indicates whether a given statement is the last one to be read before aborting the search.
Progression indicates that the assumptions made by the framework are applicable in the first place, requiring a sequential process. Decay is encoded by a diminishing attention, relating to continuation probability and weight (/), while saturation relates to the last probability (). In sum, this allows to operationalize the reading model as a monotonically decreasing weight function over statements, discounting the contribution of statements occurring later in the response. This induces a corresponding response-level document organization where the most important pieces of information come first and are then followed by increasingly insignificant details, like the inverted pyramid scheme of news articles (Pöttker, 2003).
3.4.3. Accumulation Model for Generative IR
To combine gain and discount values over all considered statements, we argue in favor of the accumulation model of expected total utility (Carterette, 2011; Moffat et al., 2013). It considers the total utility a searcher accumulates from the whole response. Alternatively, measures could be based around estimating the total ‘cost’ of accruing information from the response in terms of the effort expended (Carterette, 2011). However, we argue that, since this effort is comparatively small in text SERPs, optimizing for it is not suitable for reliably differentiating systems in evaluation.
4. Operationalizing Evaluation
A flowchart illustrating the evaluation procedure for generative ad hoc IR. It subdivides the process into the four steps of setup, statement segmentation, utility assessment, and measuring utility. These steps have further graphical elements which correspond to the process described in the text of this section.
This section considers operationalizations of the proposed user model. The goal is to take stake in what possibilities exist for each step of the process, in an effort to illustrate the required components and how they can be implemented. These considerations are summarized in Figure 6, with each component (Figure rows) described in a subsection below. The experimental setup encompasses a document collection, a set of topics reflecting the search task, and a set of generative retrieval systems to be evaluated (Section 4.1). Their responses to queries are (optionally) split into statements using a segmentation approach (Section 4.2). Statements are then assessed for their utility, distinguishing between assessment without prior reference, and assessment in relation to prior reference material (Section 4.3). Given annotations and an evaluation measure, the systems can then be ranked with respect to their effectiveness as indicated by an aggregated score (Section 4.4). In each of these four steps, we survey relevant literature and juxtapose proposed evaluation processes with regard to their advantages and disadvantages in the context of the assumed user model.
4.1. Experimental Setting
The established approach for the reproducible evaluation of traditional retrieval systems in an academic context is offline evaluation (Cleverdon, 1967; Sanderson, 2010). It encompasses a document collection, a set of topics reflecting the information needs stated by users, and the set of systems to be tested. Generative retrieval evaluation does not diverge from this basic procedure. Yet, the set of topics should include ones that reflect the search task for which generative retrieval systems are useful, i.e., the synthetic task posited in Section 2.3. Furthermore, a ranking of documents could be pre-supplied for each topic’s query in order to exclusively study the systems’ synthesizing ability. These can be taken from a baseline retrieval system, shared task results (Craswell et al., 2020, 2021, 2022), or query logs (Reimer et al., 2023). While opting for offline evaluation allows to reuse established experiment infrastructure such as the TREC format specifications for run and utility judgment files,666https://github.com/usnistgov/trec_eval/ generative retrieval systems introduce new requirements. Specifically, a run file represents a text SERP, and should thus include the generated text instead of a ranked list of document identifiers. Utility judgments should be persisted together with the annotated text, since no static document identifiers are available.
4.2. Segmenting Statements
While the complete response provided by the system can be annotated as-is (this is especially warranted for response-level utility), in order to ease annotation, it can be segmented into retrieval units (suitable for statement-level utility). This approach of subdividing a response into smaller units is well established in evaluating generated texts in NLP (Liu et al., 2023a; Nenkova et al., 2007; Dang and Lin, 2007), and has been proposed for IR as well (Sakai et al., 2011; Sakai, 2023). Unit statements should be atomic, in the sense that an assessor should be able to make an informed and reliable decision about their utility with little to no surrounding context.
To this end, human judges can be employed to extract statements (Dang et al., 2006; Dang and Lin, 2007), but the high effort and low repeatability, as well as the inability to assess the effectiveness of a new system without repeated human intervention renders this approach impractical in most settings. Automatic means of statement segmentation, comparable to the established task of web page segmentation (Kiesel et al., 2021), could include splitting after each given reference (useful for experiments investigating grounding, as each statement has a clear attributable source), sentence-level splitting (useful for fine-grained utility dimensions such as correctness or coverage), or prompting the model to output already delineated statements.
4.3. Assessing Utility
Two different settings for collecting utility assessments can be discerned: (1) a direct assessment of the responses is carried out, without comparing to a separate ground truth; and (2) the unjudged responses can be compared to pre-existing reference responses on the same document and/or query set. The first is similar to reference-free evaluation in summarization (Fabbri et al., 2021), which instructs annotators to assess the summary directly, while the second is similar to reference-based evaluation in summarization (Bhandari et al., 2020), which instructs annotators to assess the overlap between the system output and reference response, under the assumption that the reference response is the gold standard, or at least exemplary of utility. Not all utility dimensions can be judged on the generated text alone (as, e.g., clarity of language can), but also require information beyond the generated text (e.g., topical/factual correctness). We therefore discern reference responses and context: reference responses are one or more pre-existing texts to which a new response is compared, while context covers the assessment information required. An assessment made with context only is therefore deemed reference-free.
Reference-Free Assessment
To operationalize reference-free evaluation for generative IR, the straightforward approach is to task human judges with assessing a given output. Yet, possibilities also include using the self-reported uncertainty of generative models with out-of-domain data (Nalisnick et al., 2019), or relying on other generative models to assess the quality of the output, such as BARTScore (Yuan et al., 2021) or GPTScore (Fu et al., 2023). Classifiers trained to estimate the magnitude of a utility dimension have also been used (Kulesza and Shieber, 2004). Ranking, either in a pairwise or listwise fashion is an additional form of assessment, i.e., tasking a judge with ordering statements of unknown utility with respect to a given utility dimension (Gienapp et al., 2020), under the hypothesis that a response with higher utility will be ranked higher, too.
Reference-Based Assessment
To operationalize reference-based assessment, commonly a similarity measure is applied between reference and response. Lazaridou et al. (2022) evaluate their generative retrieval system for the task of question answering by matching words between generated response and the gold answer. Similarity, Arabzadeh et al. (2024) assign relevance scores to candidate answers in a QA task by measuring their similarity to annotated ground truth data in latent space. Other content overlap metrics, though not necessarily transferable to the setup proposed here, such as BLEU (Papineni et al., 2002), NIST (Doddington, 2002), ROUGE (Lin, 2004) TER (Snover et al., 2006), METEOR (Banerjee and Lavie, 2005), BERT Score (Zhang et al., 2020), or MoverScore (Zhao et al., 2019) have been used to compare a generated text to a reference text, either in full or at the statement level. Ranking models have also proven useful for the relative assessment of generated texts in comparison to references, e.g., in machine translation (Duh, 2008; Song and Cohn, 2011), both in a listwise (Li et al., 2013) as well as a pairwise setting (Guzmán et al., 2014, 2015). Arabzadeh et al. (2024) implement a kind of pseudo-relevance feedback by retrieving candidate reference documents from a corpus, using highly-ranked ones as references.
4.4. Measuring Effectiveness
For statement-level evaluation, the individual utility of statements has to be combined into an overall score for the response. Effectiveness measures for the proposed aggregation model of expected total utility take the general form (Carterette, 2011), where is the evaluation depth, or in our case, response length, is the utility of the statement at position , and is the probability of the user aborting their search immediately after position . The former is referred to as a gain function, given by the utility assessments of statements collected before. The latter as a discount function, chosen based on prior information about typical user behavior. The widely established measures of DCG and nDCG (Järvelin and Kekäläinen, 2002) used for traditional IR evaluation stem from this family of measures (Carterette, 2011) and seem suitable for generative retrieval evaluation as well. Yet, they assume a logarithmic discount function. It is currently unclear if this is an appropriate choice to model the effect of decay and saturation in the proposed reading model for generative IR. While the family of measures is thus applicable, the concrete choice of measure needs further empirical validation from user experiments.
For response-level evaluation, two choices for measuring effectiveness exist: either utility is annotated directly for a response, or it is aggregated from individual statement utility. While the latter seems counterintuitive to the response-level vs. statement-level distinction made for utility before, note that the level of granularity on which a utility dimension is defined, and the level of granularity at which annotations are collected can differ. Response-level utility may be aggregated from annotations of individual statements, or statement utility may be derived from annotations of the whole response. For example, consider the response-level utility dimension of broad coverage. It can be estimated by measuring the breadth of topics occurring over all statements, hereby annotating which topics occur in each statement. The previously motivated family of DCG-type measures can be extended to support such evaluation. For example, measure modifications similar to -nDCG (Clarke et al., 2008) that reward a diverse set of topics in a ranked list can be made for generative IR as well. Independent of how a single score is produced for each response, the final system score is aggregated over multiple topics, increasing robustness and enabling statistical testing.
4.5. Comparison with Existing Frameworks
Two other approaches for the evaluation of generative retrieval systems have been proposed recently: SWAN (Sakai, 2023) and EXAM (Sander and Dietz, 2021). The starting point of both is a text SERP response, albeit less formalized and without considering the synthetic search task it enables.
SWAN follows a similar approach as is proposed here, first establishing the notion of ‘information nuggets’, i.e., statements, that constitute the response. Then, a total of 20 categories are described, indicating how a nugget may be scored. The individual nugget scores are then averaged over the whole response. Here, too, two different levels of score categories, i.e., utility dimensions are considered. While similar, our approach and SWAN differ in three important aspects. First, we base our method on a theoretical foundation in the form of a user model, whereas SWAN is mainly motivated from a standpoint of practicability. Second, SWAN is geared towards conversational search, while we consider the ad hoc search task. And third, the utility dimensions we propose differ from SWAN due to the shift in scope: we exclude dimensions specific to conversational search (e.g., recoverability, engagingness), and also those which do not serve to operationalize evaluation for the synthetic search task specifically (such as non-toxicity, robustness to input variations, etc.). The majority of the remaining utility dimensions from SWAN can be mapped to ours.
EXAM takes a completely different approach. Instead of directly evaluating inherent qualities of the generated text, it considers the downstream effectiveness of a Q&A system that ingests the generated answer on multiple-choice questions. The hypothesis is that the correctness of its responses are correlated with the quality of the generated text it uses as input. Being an automatic evaluation method, this allows for rapid experimentation, yet exhibits three major drawbacks: it offers no fine-grained insight into the quality of the generated text, it is not grounded in a user model, and it requires a suitable Q&A system, impacting reliability and comparability, since there are no accepted standards.
In sum, our approach can be related to existing methods in terms of compatibility, complementarity, and consistency. Our approach is compatible with SWAN, as it is derived from similar assumptions, yet adding a theoretical foundation, and constructed with a different search task in mind. Our approach is complementary to EXAM, as our focus is on fine-grained, reliable, user-oriented evaluation, whereas EXAM excels for rapid, system-oriented experimentation with little overhead. Furthermore, our approach is consistent with traditional IR evaluation techniques, making only small adaptations to the utility, browsing, and aggregation models to accommodate the new search paradigm. We believe that this renders much of the work on methods and theoretical foundation for traditional IR evaluation still applicable.
5. Conclusion
Generative retrieval introduces a new paradigm for the retrieval of information. With it comes the need to measure and understand new utility dimensions that make text SERP responses from generative retrieval systems relevant to a user’s information need. In this paper, we have extrapolated a theoretical foundation for the evaluation of generative retrieval systems from traditional IR and related disciplines. First, we established that the search task of generative ad hoc retrieval goes beyond acquiring information, and instead enables the condensation of information, a process we dub the ‘synthetic search task’. Second, we proposed a new user model that accommodates this task, including evaluation objectives, utility dimensions, and a browsing model for text SERPs. Finally, we outlined how one could operationalize the evaluation of generative retrieval systems, surveying how existing evaluation approaches relate to, and could fit into the proposed methodology.
Many techniques for constructing generative retrieval systems are currently emerging, but evaluating their output is still a non-standardized and thus hardly comparable effort, lacking a theoretical motivation. We have provided our vision of a comprehensive approach for evaluating generative retrieval systems. Yet, we believe that user experiments are needed to effectively apply this theoretical motivation, and studying its reliability and validity. This requires a meta-evaluation, such as recently started by Arabzadeh and Clarke (2024), of both, existing measures and measures modified for generative IR specifically, to study how well they align with user preferences, and to study the proposed utility dimensions and their ability to reflect user satisfaction, similar to studies conducted for traditional IR (Cambazoglu et al., 2021). In addition, investigating user interactions with generative retrieval systems is warranted; for example, are user clicks on cited documents in a generated response indicative of their relevance or the user’s disbelief, or will generative retrieval make clicks superfluous?
Limitations
The evaluation process we propose in this paper is limited in two ways. First, we opted for a holistic evaluation of text SERPs, i.e., instead of evaluating the pipeline of components that constitute the generative retrieval system individually, we focus on evaluating the final response. Second, the evaluation is additionally limited to answer the question if a generative retrieval system is successful at supporting the synthetic search task. This does not consider the more general evaluation objectives that all search systems are subject to (such as bias, fairness, ethicality, or user privacy). In that sense, our considerations are specific to generative IR, disregarding the evaluation of systemic aspects of IR as a whole. This is not meant to deemphasize the importance of evaluating, e.g., bias in search results, but rather considers it to be outside the scope of this paper.
Acknowledgements.
This publication has been partially supported by the ScaDS.AI Center for Scalable Data Analytics and Artificial Intelligence, funded by the Federal Ministry of Education and Research of Germany and by the Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus, by a Research Fellowship for Harrisen Scells from the Alexander von Humboldt Foundation, and by the OpenWebSearch.eu project, funded by the European Union (GA 101070014).References
- (1)
- Agosti et al. (2014) Maristella Agosti, Norbert Fuhr, Elaine Toms, and Pertti Vakkari. 2014. Evaluation Methodologies in Information Retrieval (Dagstuhl Seminar 13441). Dagstuhl Reports 3, 10 (2014), 92–126.
- Alaofi et al. (2023) Marwah Alaofi, Luke Gallagher, Mark Sanderson, Falk Scholer, and Paul Thomas. 2023. Can Generative LLMs Create Query Variants for Test Collections? An Exploratory Study. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2023, Taipei, Taiwan, July 23–27, 2023, Hsin-Hsi Chen, Wei-Jou (Edward) Duh, Hen-Hsen Huang, Makoto P. Kato, Josiane Mothe, and Barbara Poblete (Eds.). ACM, 1869–1873.
- Alkaissi and McFarlane (2023) Hussam Alkaissi and Samy I. McFarlane. 2023. Artificial Hallucinations in ChatGPT: Implications in Scientific Writing. Cureus 15, 2 (2023), 4 pages.
- Arabzadeh et al. (2024) Negar Arabzadeh, Amin Bigdeli, and Charles L. A. Clarke. 2024. Adapting Standard Retrieval Benchmarks to Evaluate Generated Answers. In Advances in Information Retrieval - 46th European Conference on Information Retrieval, ECIR 2024, Glasgow, UK, March 24–28, 2024, Proceedings, Part II (Lecture Notes in Computer Science, Vol. 14609), Nazli Goharian, Nicola Tonellotto, Yulan He, Aldo Lipani, Graham McDonald, Craig Macdonald, and Iadh Ounis (Eds.). Springer, 399–414.
- Arabzadeh and Clarke (2024) Negar Arabzadeh and Charles L. A. Clarke. 2024. A Comparison of Methods for Evaluating Generative IR. arXiv 2404.04044.
- Arora et al. (2023) Daman Arora, Anush Kini, Sayak Ray Chowdhury, Nagarajan Natarajan, Gaurav Sinha, and Amit Sharma. 2023. GAR-meets-RAG Paradigm for Zero-Shot Information Retrieval. arXiv 2310.20158.
- AutoGPT Contributors (2023) AutoGPT Contributors. 2023. AutoGPT: The Heart of the Open-Source Agent Ecosystem. https://github.com/Significant-Gravitas/AutoGPT.
- Banerjee and Lavie (2005) Satanjeev Banerjee and Alon Lavie. 2005. METEOR: An Automatic Metric for MT Evaluation with Improved Correlation with Human Judgments. In Proceedings of the Workshop on Intrinsic and Extrinsic Evaluation Measures for Machine Translation and/or Summarization@ACL 2005, Ann Arbor, Michigan, USA, June 29, 2005, Jade Goldstein, Alon Lavie, Chin-Yew Lin, and Clare R. Voss (Eds.). Association for Computational Linguistics, 65–72.
- Bauer et al. (2023) Christine Bauer, Ben Carterette, Nicola Ferro, and Norbert Fuhr. 2023. Report from Dagstuhl Seminar 23031: Frontiers of Information Access Experimentation for Research and Education. arXiv 2305.01509.
- Bénédict et al. (2023) Gabriel Bénédict, Ruqing Zhang, and Donald Metzler. 2023. Gen-IR @ SIGIR 2023: The First Workshop on Generative Information Retrieval. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2023, Taipei, Taiwan, July 23-27, 2023, Hsin-Hsi Chen, Wei-Jou (Edward) Duh, Hen-Hsen Huang, Makoto P. Kato, Josiane Mothe, and Barbara Poblete (Eds.). ACM, 3460–3463. https://doi.org/10.1145/3539618.3591923
- Bevilacqua et al. (2022) Michele Bevilacqua, Giuseppe Ottaviano, Patrick S. H. Lewis, Scott Yih, Sebastian Riedel, and Fabio Petroni. 2022. Autoregressive Search Engines: Generating Substrings as Document Identifiers. In Advances in Neural Information Processing Systems 35: Annual Conference on Neural Information Processing Systems 2022, NeurIPS 2022, New Orleans, LA, USA, November 28 – December 9, 2022. 16 pages.
- Bhandari et al. (2020) Manik Bhandari, Pranav Narayan Gour, Atabak Ashfaq, Pengfei Liu, and Graham Neubig. 2020. Re-evaluating Evaluation in Text Summarization. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing, EMNLP 2020, Online, November 16–20, 2020, Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, 9347–9359.
- Borgeaud et al. (2022) Sebastian Borgeaud, Arthur Mensch, Jordan Hoffmann, Trevor Cai, Eliza Rutherford, Katie Millican, George van den Driessche, Jean-Baptiste Lespiau, Bogdan Damoc, Aidan Clark, Diego de Las Casas, Aurelia Guy, Jacob Menick, Roman Ring, Tom Hennigan, Saffron Huang, Loren Maggiore, Chris Jones, Albin Cassirer, Andy Brock, Michela Paganini, Geoffrey Irving, Oriol Vinyals, Simon Osindero, Karen Simonyan, Jack W. Rae, Erich Elsen, and Laurent Sifre. 2022. Improving Language Models by Retrieving from Trillions of Tokens. In International Conference on Machine Learning, ICML 2022, 17–23 July 2022, Baltimore, Maryland, USA (Proceedings of Machine Learning Research, Vol. 162), Kamalika Chaudhuri, Stefanie Jegelka, Le Song, Csaba Szepesvári, Gang Niu, and Sivan Sabato (Eds.). PMLR, 2206–2240.
- Broder (2002) Andrei Z. Broder. 2002. A Taxonomy of Web Search. SIGIR Forum 36, 2 (2002), 3–10.
- Buscher et al. (2012) Georg Buscher, Andreas Dengel, Ralf Biedert, and Ludger van Elst. 2012. Attentive Documents: Eye Tracking as Implicit Feedback for Information Retrieval and Beyond. ACM Trans. Interact. Intell. Syst. 1, 2 (2012), 9:1–9:30.
- Câmara et al. (2022) Arthur Câmara, David Maxwell, and Claudia Hauff. 2022. Searching, Learning, and Subtopic Ordering: A Simulation-Based Analysis. In Advances in Information Retrieval - 44th European Conference on IR Research, ECIR 2022, Stavanger, Norway, April 10–14, 2022, Proceedings, Part I (Lecture Notes in Computer Science, Vol. 13185), Matthias Hagen, Suzan Verberne, Craig Macdonald, Christin Seifert, Krisztian Balog, Kjetil Nørvåg, and Vinay Setty (Eds.). Springer, 142–156.
- Cambazoglu et al. (2021) Berkant Barla Cambazoglu, Valeria Bolotova-Baranova, Falk Scholer, Mark Sanderson, Leila Tavakoli, and W. Bruce Croft. 2021. Quantifying Human-Perceived Answer Utility in Non-factoid Question Answering. In CHIIR ’21: ACM SIGIR Conference on Human Information Interaction and Retrieval, Canberra, ACT, Australia, March 14–19, 2021, Falk Scholer, Paul Thomas, David Elsweiler, Hideo Joho, Noriko Kando, and Catherine Smith (Eds.). ACM, 75–84.
- Capra and Arguello (2023) Robert Capra and Jaime Arguello. 2023. How does AI Chat Change Search Behaviors? arXiv 2307.03826.
- Carterette (2011) Ben Carterette. 2011. System Effectiveness, User Models, and User Utility: A Conceptual Framework for Investigation. In Proceeding of the 34th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2011, Beijing, China, July 25–29, 2011, Wei-Ying Ma, Jian-Yun Nie, Ricardo Baeza-Yates, Tat-Seng Chua, and W. Bruce Croft (Eds.). ACM, 903–912.
- Chen et al. (2022) Jiangui Chen, Ruqing Zhang, Jiafeng Guo, Yixing Fan, and Xueqi Cheng. 2022. GERE: Generative Evidence Retrieval for Fact Verification. In SIGIR ’22: The 45th International ACM SIGIR Conference on Research and Development in Information Retrieval, Madrid, Spain, July 11 – 15, 2022, Enrique Amigó, Pablo Castells, Julio Gonzalo, Ben Carterette, J. Shane Culpepper, and Gabriella Kazai (Eds.). ACM, 2184–2189.
- Chen et al. (2023) Tong Chen, Hongwei Wang, Sihao Chen, Wenhao Yu, Kaixin Ma, Xinran Zhao, Hongming Zhang, and Dong Yu. 2023. Dense X Retrieval: What Retrieval Granularity Should We Use? arXiv 2312.06648.
- Clarke et al. (2008) Charles L. A. Clarke, Maheedhar Kolla, Gordon V. Cormack, Olga Vechtomova, Azin Ashkan, Stefan Büttcher, and Ian MacKinnon. 2008. Novelty and Diversity in Information Retrieval Evaluation. In Proceedings of the 31st Annual International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2008, Singapore, July 20–24, 2008, Sung-Hyon Myaeng, Douglas W. Oard, Fabrizio Sebastiani, Tat-Seng Chua, and Mun-Kew Leong (Eds.). ACM, 659–666.
- Cleverdon (1967) Cyril W. Cleverdon. 1967. The Cranfield Tests on Index Language Devices. Aslib Proceedings 19, 6 (1967), 173–194.
- Craswell et al. (2020) Nick Craswell, Bhaskar Mitra, Emine Yilmaz, and Daniel Campos. 2020. Overview of the TREC 2020 Deep Learning Track. In Proceedings of the Twenty-Ninth Text REtrieval Conference, TREC 2020, Virtual Event [Gaithersburg, Maryland, USA], November 16–20, 2020 (NIST Special Publication, Vol. 1266), Ellen M. Voorhees and Angela Ellis (Eds.). National Institute of Standards and Technology (NIST), 13 pages.
- Craswell et al. (2021) Nick Craswell, Bhaskar Mitra, Emine Yilmaz, Daniel Campos, and Jimmy Lin. 2021. Overview of the TREC 2021 Deep Learning Track. In Proceedings of the Thirtieth Text REtrieval Conference, TREC 2021, online, November 15–19, 2021 (NIST Special Publication, Vol. 500-335), Ian Soboroff and Angela Ellis (Eds.). National Institute of Standards and Technology (NIST), 16 pages.
- Craswell et al. (2022) Nick Craswell, Bhaskar Mitra, Emine Yilmaz, Daniel Campos, Jimmy Lin, Ellen M. Voorhees, and Ian Soboroff. 2022. Overview of the TREC 2022 Deep Learning Track. In Proceedings of the Thirty-First Text REtrieval Conference, TREC 2022, online, November 15–19, 2022 (NIST Special Publication, Vol. 500-338), Ian Soboroff and Angela Ellis (Eds.). National Institute of Standards and Technology (NIST), 21 pages.
- Culpepper et al. (2018) J. Shane Culpepper, Fernando Diaz, and Mark D. Smucker. 2018. Research Frontiers in Information Retrieval: Report from the Third Strategic Workshop on Information Retrieval in Lorne (SWIRL 2018). SIGIR Forum 52, 1 (2018), 34–90.
- Dang (2005) Hoa Trang Dang. 2005. Overview of DUC 2005. In DUC 2005, Document Understanding Workshop October 9–10, 2005, Vancouver, B.C., Canada. 1–12.
- Dang and Lin (2007) Hoa Trang Dang and Jimmy Lin. 2007. Different Structures for Evaluating Answers to Complex Questions: Pyramids Won’t Topple, and Neither Will Human Assessors. In ACL 2007, Proceedings of the 45th Annual Meeting of the Association for Computational Linguistics, June 23–30, 2007, Prague, Czech Republic, John Carroll, Antal van den Bosch, and Annie Zaenen (Eds.). The Association for Computational Linguistics, 768–775.
- Dang et al. (2006) Hoa Trang Dang, Jimmy Lin, and Diane Kelly. 2006. Overview of the TREC 2006 Question Answering Track. In Proceedings of the Fifteenth Text REtrieval Conference, TREC 2006, Gaithersburg, Maryland, USA, November 14–17, 2006 (NIST Special Publication, Vol. 500-272), Ellen M. Voorhees and Lori P. Buckland (Eds.). National Institute of Standards and Technology (NIST), 18 pages.
- De Cao et al. (2021) Nicola De Cao, Gautier Izacard, Sebastian Riedel, and Fabio Petroni. 2021. Autoregressive Entity Retrieval. In 9th International Conference on Learning Representations, ICLR 2021, Virtual Event, Austria, May 3–7, 2021. OpenReview.net, 20 pages.
- Deckers et al. (2023) Niklas Deckers, Maik Fröbe, Johannes Kiesel, Gianluca Pandolfo, Christopher Schröder, Benno Stein, and Martin Potthast. 2023. The Infinite Index: Information Retrieval on Generative Text-To-Image Models. In Proceedings of the 2023 Conference on Human Information Interaction and Retrieval, CHIIR 2023, Austin, TX, USA, March 19–23, 2023, Jacek Gwizdka and Soo Young Rieh (Eds.). ACM, 172–186.
- Doddington (2002) George R. Doddington. 2002. Automatic Evaluation of Machine Translation Quality Using N-gram Co-occurrence Statistics. In Proceedings of the Second International Conference on Human Language Technology Research, HLT 2002. 138–145.
- Duh (2008) Kevin Duh. 2008. Ranking vs. Regression in Machine Translation Evaluation. In Proceedings of the Third Workshop on Statistical Machine Translation, WMT@ACL 2008, Columbus, Ohio, USA, June 19, 2008, Chris Callison-Burch, Philipp Koehn, Christof Monz, Josh Schroeder, and Cameron S. Fordyce (Eds.). Association for Computational Linguistics, 191–194.
- Fabbri et al. (2021) Alexander R. Fabbri, Wojciech Kryscinski, Bryan McCann, Caiming Xiong, Richard Socher, and Dragomir R. Radev. 2021. SummEval: Re-evaluating Summarization Evaluation. Trans. Assoc. Comput. Linguistics 9 (2021), 391–409.
- Faggioli et al. (2023) Guglielmo Faggioli, Laura Dietz, Charles L. A. Clarke, Gianluca Demartini, Matthias Hagen, Claudia Hauff, Noriko Kando, Evangelos Kanoulas, Martin Potthast, Benno Stein, and Henning Wachsmuth. 2023. Perspectives on Large Language Models for Relevance Judgment. In Proceedings of the 2023 ACM SIGIR International Conference on Theory of Information Retrieval, ICTIR 2023, Taipei, Taiwan, 23 July 2023, Masaharu Yoshioka, Julia Kiseleva, and Mohammad Aliannejadi (Eds.). ACM, 39–50.
- Formal et al. (2021) Thibault Formal, Carlos Lassance, Benjamin Piwowarski, and Stéphane Clinchant. 2021. SPLADE v2: Sparse Lexical and Expansion Model for Information Retrieval. arXiv 2109.10086.
- Frey et al. (2013) Aline Frey, Gelu Ionescu, Benoit Lemaire, Francisco López-Orozco, Thierry Baccino, and Anne Guérin-Dugué. 2013. Decision-Making in Information Seeking on Texts: An Eye-Fixation-Related Potentials Investigation. Frontiers in Systems Neuroscience 7 (2013), 22 pages.
- Fröbe et al. (2023) Maik Fröbe, Lukas Gienapp, Martin Potthast, and Matthias Hagen. 2023. Bootstrapped nDCG Estimation in the Presence of Unjudged Documents. In Advances in Information Retrieval - 45th European Conference on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2–6, 2023, Proceedings, Part I (Lecture Notes in Computer Science, Vol. 13980), Jaap Kamps, Lorraine Goeuriot, Fabio Crestani, Maria Maistro, Hideo Joho, Brian Davis, Cathal Gurrin, Udo Kruschwitz, and Annalina Caputo (Eds.). Springer, 313–329.
- Fu et al. (2023) Jinlan Fu, See-Kiong Ng, Zhengbao Jiang, and Pengfei Liu. 2023. GPTScore: Evaluate as You Desire. arXiv 2302.04166.
- Gao et al. (2021) Tianyu Gao, Adam Fisch, and Danqi Chen. 2021. Making Pre-trained Language Models Better Few-shot Learners. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing, ACL/IJCNLP 2021, (Volume 1: Long Papers), Virtual Event, August 1–6, 2021, Chengqing Zong, Fei Xia, Wenjie Li, and Roberto Navigli (Eds.). Association for Computational Linguistics, 3816–3830.
- Gao et al. (2023) Yunfan Gao, Yun Xiong, Xinyu Gao, Kangxiang Jia, Jinliu Pan, Yuxi Bi, Yi Dai, Jiawei Sun, Qianyu Guo, Meng Wang, and Haofen Wang. 2023. Retrieval-Augmented Generation for Large Language Models: A Survey. arXiv 2312.10997.
- Gienapp et al. (2020) Lukas Gienapp, Benno Stein, Matthias Hagen, and Martin Potthast. 2020. Efficient Pairwise Annotation of Argument Quality. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, ACL 2020, Online, July 5–10, 2020, Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel R. Tetreault (Eds.). Association for Computational Linguistics, 5772–5781.
- Gospodinov et al. (2023) Mitko Gospodinov, Sean MacAvaney, and Craig Macdonald. 2023. Doc2Query-: When Less is More. In Advances in Information Retrieval - 45th European Conference on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2–6, 2023, Proceedings, Part II (Lecture Notes in Computer Science, Vol. 13981), Jaap Kamps, Lorraine Goeuriot, Fabio Crestani, Maria Maistro, Hideo Joho, Brian Davis, Cathal Gurrin, Udo Kruschwitz, and Annalina Caputo (Eds.). Springer, 414–422.
- Guu et al. (2020) Kelvin Guu, Kenton Lee, Zora Tung, Panupong Pasupat, and Ming-Wei Chang. 2020. REALM: Retrieval-Augmented Language Model Pre-Training. arXiv 2002.08909.
- Guzmán et al. (2014) Francisco Guzmán, Shafiq R. Joty, Lluís Màrquez, Alessandro Moschitti, Preslav Nakov, and Massimo Nicosia. 2014. Learning to Differentiate Better from Worse Translations. In Proceedings of the 2014 Conference on Empirical Methods in Natural Language Processing, EMNLP 2014, October 25–29, 2014, Doha, Qatar, A meeting of SIGDAT, a Special Interest Group of the ACL, Alessandro Moschitti, Bo Pang, and Walter Daelemans (Eds.). ACL, 214–220.
- Guzmán et al. (2015) Francisco Guzmán, Shafiq R. Joty, Lluís Màrquez, and Preslav Nakov. 2015. Pairwise Neural Machine Translation Evaluation. In Proceedings of the 53rd Annual Meeting of the Association for Computational Linguistics and the 7th International Joint Conference on Natural Language Processing of the Asian Federation of Natural Language Processing, ACL 2015, July 26–31, 2015, Beijing, China, Volume 1: Long Papers. The Association for Computer Linguistics, 805–814.
- Gwizdka (2014) Jacek Gwizdka. 2014. Characterizing Relevance with Eye-Tracking Measures. In Fifth Information Interaction in Context Symposium, IIiX ’14, Regensburg, Germany, August 26–29, 2014, David Elsweiler, Bernd Ludwig, Leif Azzopardi, and Max L. Wilson (Eds.). ACM, 58–67.
- Hearst (2009) Marti A. Hearst. 2009. Search User Interfaces. Cambridge University Press.
- Huang et al. (2021) Yi-Chong Huang, Xia-Chong Feng, Xiao-Cheng Feng, and Bing Qin. 2021. The Factual Inconsistency Problem in Abstractive Text Summarization: A Survey. arXiv 2104.14839.
- Izacard and Grave (2021) Gautier Izacard and Edouard Grave. 2021. Leveraging Passage Retrieval with Generative Models for Open Domain Question Answering. In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Main Volume, EACL 2021, Online, April 19–23, 2021, Paola Merlo, Jörg Tiedemann, and Reut Tsarfaty (Eds.). Association for Computational Linguistics, 874–880.
- Järvelin and Kekäläinen (2002) Kalervo Järvelin and Jaana Kekäläinen. 2002. Cumulated Gain-Based Evaluation of IR Techniques. ACM Trans. Inf. Syst. 20, 4 (2002), 422–446.
- Ji et al. (2023) Ziwei Ji, Nayeon Lee, Rita Frieske, Tiezheng Yu, Dan Su, Yan Xu, Etsuko Ishii, Yejin Bang, Andrea Madotto, and Pascale Fung. 2023. Survey of Hallucination in Natural Language Generation. ACM Comput. Surv. 55, 12 (2023), 248:1–248:38.
- Jiang et al. (2022) Zhengbao Jiang, Luyu Gao, Zhiruo Wang, Jun Araki, Haibo Ding, Jamie Callan, and Graham Neubig. 2022. Retrieval as Attention: End-to-end Learning of Retrieval and Reading within a Single Transformer. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, EMNLP 2022, Abu Dhabi, United Arab Emirates, December 7-11, 2022, Yoav Goldberg, Zornitsa Kozareva, and Yue Zhang (Eds.). Association for Computational Linguistics, 2336–2349.
- Jiang et al. (2023) Zhengbao Jiang, Frank F. Xu, Luyu Gao, Zhiqing Sun, Qian Liu, Jane Dwivedi-Yu, Yiming Yang, Jamie Callan, and Graham Neubig. 2023. Active Retrieval Augmented Generation. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, EMNLP 2023, Singapore, December 6-10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, 7969–7992.
- Jin et al. (2020) Di Jin, Zhijing Jin, Joey Tianyi Zhou, Lisa Orii, and Peter Szolovits. 2020. Hooks in the Headline: Learning to Generate Headlines with Controlled Styles. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, ACL 2020, Online, July 5–10, 2020, Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel R. Tetreault (Eds.). Association for Computational Linguistics, 5082–5093.
- Kelly (2009) Diane Kelly. 2009. Methods for Evaluating Interactive Information Retrieval Systems with Users. Found. Trends Inf. Retr. 3, 1-2 (2009), 1–224.
- Khattab et al. (2022) Omar Khattab, Keshav Santhanam, Xiang Lisa Li, David Hall, Percy Liang, Christopher Potts, and Matei Zaharia. 2022. Demonstrate-Search-Predict: Composing Retrieval and Language Models for Knowledge-Intensive NLP. arXiv 2212.14024.
- Khot et al. (2023) Tushar Khot, Harsh Trivedi, Matthew Finlayson, Yao Fu, Kyle Richardson, Peter Clark, and Ashish Sabharwal. 2023. Decomposed Prompting: A Modular Approach for Solving Complex Tasks. In The Eleventh International Conference on Learning Representations, ICLR 2023, Kigali, Rwanda, May 1–5, 2023. OpenReview.net, 69 pages.
- Kiciman et al. (2023) Emre Kiciman, Robert Ness, Amit Sharma, and Chenhao Tan. 2023. Causal Reasoning and Large Language Models: Opening a New Frontier for Causality. arXiv 2305.00050.
- Kiesel et al. (2021) Johannes Kiesel, Lars Meyer, Florian Kneist, Benno Stein, and Martin Potthast. 2021. An Empirical Comparison of Web Page Segmentation Algorithms. In Advances in Information Retrieval - 43rd European Conference on IR Research, ECIR 2021, Virtual Event, March 28 – April 1, 2021, Proceedings, Part II (Lecture Notes in Computer Science, Vol. 12657), Djoerd Hiemstra, Marie-Francine Moens, Josiane Mothe, Raffaele Perego, Martin Potthast, and Fabrizio Sebastiani (Eds.). Springer, 62–74.
- Koopman et al. (2023) Bevan Koopman, Ahmed Mourad, Hang Li, Anton van der Vegt, Shengyao Zhuang, Simon Gibson, Yash Dang, David Lawrence, and Guido Zuccon. 2023. AgAsk: An Agent to Help Answer Farmer’s Questions from Scientific Documents. International Journal on Digital Libraries (2023), 16 pages.
- Koopman and Zuccon (2023) Bevan Koopman and Guido Zuccon. 2023. Dr ChatGPT Tell Me What I Want to Hear: How Different Prompts Impact Health Answer Correctness. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, EMNLP 2023, Singapore, December 6–10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, 15012–15022. https://doi.org/10.18653/V1/2023.EMNLP-MAIN.928
- Kulesza and Shieber (2004) Alex Kulesza and Stuart M. Shieber. 2004. A Learning Approach to Improving Sentence-Level MT Evaluation. In Proceedings of the 10th Conference on Theoretical and Methodological Issues in Machine Translation of Natural Languages. 10 pages.
- Lazaridou et al. (2022) Angeliki Lazaridou, Elena Gribovskaya, Wojciech Stokowiec, and Nikolai Grigorev. 2022. Internet-Augmented Language Models Through Few-Shot Prompting for Open-Domain Question Answering. arXiv 2203.05115.
- Lewis et al. (2020) Patrick S. H. Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, Sebastian Riedel, and Douwe Kiela. 2020. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. In Advances in Neural Information Processing Systems 33: Annual Conference on Neural Information Processing Systems 2020, NeurIPS 2020, December 6–12, 2020, virtual, Hugo Larochelle, Marc’Aurelio Ranzato, Raia Hadsell, Maria-Florina Balcan, and Hsuan-Tien Lin (Eds.). 16 pages.
- Li et al. (2013) Maoxi Li, Aiwen Jiang, and Mingwen Wang. 2013. Listwise Approach to Learning to Rank for Automatic Evaluation of Machine Translation. In Proceedings of Machine Translation Summit XIV: Papers, MTSummit 2013, Nice, France, September 2–6, 2013, Andy Way, Khalil Sima’an, and Mikel L. Forcada (Eds.). 8 pages.
- Li et al. (2018) Xiangsheng Li, Yiqun Liu, Jiaxin Mao, Zexue He, Min Zhang, and Shaoping Ma. 2018. Understanding Reading Attention Distribution during Relevance Judgement. In Proceedings of the 27th ACM International Conference on Information and Knowledge Management, CIKM 2018, Torino, Italy, October 22–26, 2018, Alfredo Cuzzocrea, James Allan, Norman W. Paton, Divesh Srivastava, Rakesh Agrawal, Andrei Z. Broder, Mohammed J. Zaki, K. Selçuk Candan, Alexandros Labrinidis, Assaf Schuster, and Haixun Wang (Eds.). ACM, 733–742.
- Li et al. (2019) Xiangsheng Li, Jiaxin Mao, Chao Wang, Yiqun Liu, Min Zhang, and Shaoping Ma. 2019. Teach Machine How to Read: Reading Behavior Inspired Relevance Estimation. In Proceedings of the 42nd International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2019, Paris, France, July 21–25, 2019, Benjamin Piwowarski, Max Chevalier, Éric Gaussier, Yoelle Maarek, Jian-Yun Nie, and Falk Scholer (Eds.). ACM, 795–804.
- Lin (2004) Chin-Yew Lin. 2004. ROUGE: A Package for Automatic Evaluation of Summaries. In Text Summarization Branches Out. Association for Computational Linguistics, 74–81.
- Liu et al. (2021) Chang Liu, Ying-Hsang Liu, Jingjing Liu, and Ralf Bierig. 2021. Search Interface Design and Evaluation. Found. Trends Inf. Retr. 15, 3-4 (2021), 243–416.
- Liu et al. (2023b) Nelson F. Liu, Tianyi Zhang, and Percy Liang. 2023b. Evaluating Verifiability in Generative Search Engines. In Findings of the Association for Computational Linguistics: EMNLP 2023, Singapore, December 6-10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, 7001–7025.
- Liu and Chilton (2022) Vivian Liu and Lydia B. Chilton. 2022. Design Guidelines for Prompt Engineering Text-to-Image Generative Models. In CHI ’22: CHI Conference on Human Factors in Computing Systems, New Orleans, LA, USA, 29 April 2022 – 5 May 2022, Simone D. J. Barbosa, Cliff Lampe, Caroline Appert, David A. Shamma, Steven Mark Drucker, Julie R. Williamson, and Koji Yatani (Eds.). ACM, 384:1–384:23.
- Liu et al. (2023a) Yixin Liu, Alexander R. Fabbri, Pengfei Liu, Yilun Zhao, Linyong Nan, Ruilin Han, Simeng Han, Shafiq Joty, Chien-Sheng Wu, Caiming Xiong, and Dragomir Radev. 2023a. Revisiting the Gold Standard: Grounding Summarization Evaluation with Robust Human Evaluation. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), ACL 2023, Toronto, Canada, July 9–14, 2023, Anna Rogers, Jordan L. Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, 4140–4170.
- Lux et al. (2020) Klaus-Michael Lux, Maya Sappelli, and Martha A. Larson. 2020. Truth or Error? Towards Systematic Analysis of Factual Errors in Abstractive Summaries. In Proceedings of the First Workshop on Evaluation and Comparison of NLP Systems, Eval4NLP 2020, Online, November 20, 2020, Steffen Eger, Yang Gao, Maxime Peyrard, Wei Zhao, and Eduard H. Hovy (Eds.). Association for Computational Linguistics, 1–10.
- Ma et al. (2023) Xueguang Ma, Xinyu Zhang, Ronak Pradeep, and Jimmy Lin. 2023. Zero-Shot Listwise Document Reranking with a Large Language Model. arXiv 2305.02156.
- MacAvaney et al. (2021) Sean MacAvaney, Craig Macdonald, Roderick Murray-Smith, and Iadh Ounis. 2021. IntenT5: Search Result Diversification using Causal Language Models. arXiv 2108.04026.
- MacAvaney et al. (2020) Sean MacAvaney, Franco Maria Nardini, Raffaele Perego, Nicola Tonellotto, Nazli Goharian, and Ophir Frieder. 2020. Expansion via Prediction of Importance with Contextualization. In Proceedings of the 43rd International ACM SIGIR conference on research and development in Information Retrieval, SIGIR 2020, Virtual Event, China, July 25–30, 2020, Jimmy X. Huang, Yi Chang, Xueqi Cheng, Jaap Kamps, Vanessa Murdock, Ji-Rong Wen, and Yiqun Liu (Eds.). ACM, 1573–1576.
- Maddalena et al. (2017) Eddy Maddalena, Stefano Mizzaro, Falk Scholer, and Andrew Turpin. 2017. On Crowdsourcing Relevance Magnitudes for Information Retrieval Evaluation. ACM Trans. Inf. Syst. 35, 3 (2017), 19:1–19:32.
- Manning et al. (2008) Christopher D. Manning, Prabhakar Raghavan, and Hinrich Schütze. 2008. Introduction to Information Retrieval. Cambridge University Press.
- Maxwell and Azzopardi (2016) David Maxwell and Leif Azzopardi. 2016. Agents, Simulated Users and Humans: An Analysis of Performance and Behaviour. In Proceedings of the 25th ACM International Conference on Information and Knowledge Management, CIKM 2016, Indianapolis, IN, USA, October 24–28, 2016, Snehasis Mukhopadhyay, ChengXiang Zhai, Elisa Bertino, Fabio Crestani, Javed Mostafa, Jie Tang, Luo Si, Xiaofang Zhou, Yi Chang, Yunyao Li, and Parikshit Sondhi (Eds.). ACM, 731–740.
- Maxwell et al. (2015) David Maxwell, Leif Azzopardi, Kalervo Järvelin, and Heikki Keskustalo. 2015. Searching and Stopping: An Analysis of Stopping Rules and Strategies. In Proceedings of the 24th ACM International Conference on Information and Knowledge Management, CIKM 2015, Melbourne, VIC, Australia, October 19–23, 2015, James Bailey, Alistair Moffat, Charu C. Aggarwal, Maarten de Rijke, Ravi Kumar, Vanessa Murdock, Timos K. Sellis, and Jeffrey Xu Yu (Eds.). ACM, 313–322.
- Maxwell et al. (2017) David Maxwell, Leif Azzopardi, and Yashar Moshfeghi. 2017. A Study of Snippet Length and Informativeness: Behaviour, Performance and User Experience. In Proceedings of the 40th International ACM SIGIR Conference on Research and Development in Information Retrieval, Shinjuku, Tokyo, Japan, August 7–11, 2017, Noriko Kando, Tetsuya Sakai, Hideo Joho, Hang Li, Arjen P. de Vries, and Ryen W. White (Eds.). ACM, 135–144.
- Maynez et al. (2020) Joshua Maynez, Shashi Narayan, Bernd Bohnet, and Ryan T. McDonald. 2020. On Faithfulness and Factuality in Abstractive Summarization. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, ACL 2020, Online, July 5–10, 2020, Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel R. Tetreault (Eds.). Association for Computational Linguistics, 1906–1919.
- Mialon et al. (2023) Grégoire Mialon, Roberto Dessì, Maria Lomeli, Christoforos Nalmpantis, Ramakanth Pasunuru, Roberta Raileanu, Baptiste Rozière, Timo Schick, Jane Dwivedi-Yu, Asli Celikyilmaz, Edouard Grave, Yann LeCun, and Thomas Scialom. 2023. Augmented Language Models: A Survey. arXiv 2302.07842.
- Miller et al. (2017) Alexander H. Miller, Will Feng, Dhruv Batra, Antoine Bordes, Adam Fisch, Jiasen Lu, Devi Parikh, and Jason Weston. 2017. ParlAI: A Dialog Research Software Platform. In Proceedings of the 2017 Conference on Empirical Methods in Natural Language Processing, EMNLP 2017, Copenhagen, Denmark, September 9–11, 2017 - System Demonstrations, Lucia Specia, Matt Post, and Michael Paul (Eds.). Association for Computational Linguistics, 79–84.
- Moffat et al. (2015) Alistair Moffat, Peter Bailey, Falk Scholer, and Paul Thomas. 2015. INST: An Adaptive Metric for Information Retrieval Evaluation. In Proceedings of the 20th Australasian Document Computing Symposium, ADCS 2015, Parramatta, NSW, Australia, December 8–9, 2015, Laurence Anthony F. Park and Sarvnaz Karimi (Eds.). ACM, 5:1–5:4.
- Moffat et al. (2017) Alistair Moffat, Peter Bailey, Falk Scholer, and Paul Thomas. 2017. Incorporating User Expectations and Behavior into the Measurement of Search Effectiveness. ACM Trans. Inf. Syst. 35, 3 (2017), 24:1–24:38.
- Moffat et al. (2013) Alistair Moffat, Paul Thomas, and Falk Scholer. 2013. Users Versus Models: What Observation Tells us about Effectiveness Metrics. In 22nd ACM International Conference on Information and Knowledge Management, CIKM’13, San Francisco, CA, USA, October 27 – November 1, 2013, Qi He, Arun Iyengar, Wolfgang Nejdl, Jian Pei, and Rajeev Rastogi (Eds.). ACM, 659–668.
- Nakov et al. (2021) Preslav Nakov, Giovanni Da San Martino, Tamer Elsayed, Alberto Barrón-Cedeño, Rubén Míguez, Shaden Shaar, Firoj Alam, Fatima Haouari, Maram Hasanain, Watheq Mansour, Bayan Hamdan, Zien Sheikh Ali, Nikolay Babulkov, Alex Nikolov, Gautam Kishore Shahi, Julia Maria Struß, Thomas Mandl, Mücahid Kutlu, and Yavuz Selim Kartal. 2021. Overview of the CLEF-2021 CheckThat! Lab on Detecting Check-Worthy Claims, Previously Fact-Checked Claims, and Fake News. In Experimental IR Meets Multilinguality, Multimodality, and Interaction - 12th International Conference of the CLEF Association, CLEF 2021, Virtual Event, September 21–24, 2021, Proceedings (Lecture Notes in Computer Science, Vol. 12880), K. Selçuk Candan, Bogdan Ionescu, Lorraine Goeuriot, Birger Larsen, Henning Müller, Alexis Joly, Maria Maistro, Florina Piroi, Guglielmo Faggioli, and Nicola Ferro (Eds.). Springer, 264–291.
- Nalisnick et al. (2019) Eric T. Nalisnick, Akihiro Matsukawa, Yee Whye Teh, Dilan Görür, and Balaji Lakshminarayanan. 2019. Do Deep Generative Models Know What They Don’t Know?. In 7th International Conference on Learning Representations, ICLR 2019, New Orleans, LA, USA, May 6–9, 2019. OpenReview.net, 19 pages.
- Nenkova et al. (2007) Ani Nenkova, Rebecca J. Passonneau, and Kathleen R. McKeown. 2007. The Pyramid Method: Incorporating Human Content Selection Variation in Summarization Evaluation. ACM Trans. Speech Lang. Process. 4, 2 (2007), 4.
- Nishino et al. (2019) Toru Nishino, Shotaro Misawa, Ryuji Kano, Tomoki Taniguchi, Yasuhide Miura, and Tomoko Ohkuma. 2019. Keeping Consistency of Sentence Generation and Document Classification with Multi-Task Learning. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing, EMNLP-IJCNLP 2019, Hong Kong, China, November 3–7, 2019, Kentaro Inui, Jing Jiang, Vincent Ng, and Xiaojun Wan (Eds.). Association for Computational Linguistics, 3193–3203.
- Nogueira et al. (2020) Rodrigo Frassetto Nogueira, Zhiying Jiang, Ronak Pradeep, and Jimmy Lin. 2020. Document Ranking with a Pretrained Sequence-to-Sequence Model. In Findings of the Association for Computational Linguistics: EMNLP 2020, Online Event, 16–20 November 2020 (Findings of ACL, Vol. EMNLP 2020), Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, 708–718.
- Nogueira et al. (2019) Rodrigo Frassetto Nogueira, Wei Yang, Jimmy Lin, and Kyunghyun Cho. 2019. Document Expansion by Query Prediction. arXiv 1904.08375.
- Nourani et al. (2019) Mahsan Nourani, Samia Kabir, Sina Mohseni, and Eric D. Ragan. 2019. The Effects of Meaningful and Meaningless Explanations on Trust and Perceived System Accuracy in Intelligent Systems. In Proceedings of the Seventh AAAI Conference on Human Computation and Crowdsourcing, HCOMP 2019, Stevenson, WA, USA, October 28–30, 2019, Edith Law and Jennifer Wortman Vaughan (Eds.). AAAI Press, 97–105.
- Papineni et al. (2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002. BLEU: A Method for Automatic Evaluation of Machine Translation. In Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics, July 6–12, 2002, Philadelphia, PA, USA. ACL, 311–318.
- Patwardhan and Yang (2023) Quinn Patwardhan and Grace Hui Yang. 2023. Sequencing Matters: A Generate-Retrieve-Generate Model for Building Conversational Agents. arXiv 2311.09513.
- Pöttker (2003) Horst Pöttker. 2003. News and its Communicative Quality: The Inverted Pyramid – When and Why did it Appear? Journalism Studies 4, 4 (2003), 501–511.
- Qin et al. (2023) Zhen Qin, Rolf Jagerman, Kai Hui, Honglei Zhuang, Junru Wu, Jiaming Shen, Tianqi Liu, Jialu Liu, Donald Metzler, Xuanhui Wang, and Michael Bendersky. 2023. Large Language Models are Effective Text Rankers with Pairwise Ranking Prompting. arXiv 2306.17563.
- Radev and McKeown (1998) Dragomir R. Radev and Kathleen R. McKeown. 1998. Generating Natural Language Summaries from Multiple On-Line Sources. Comput. Linguistics 24, 3 (1998), 469–500.
- Radlinski and Craswell (2017) Filip Radlinski and Nick Craswell. 2017. A Theoretical Framework for Conversational Search. In Proceedings of the 2017 Conference on Conference Human Information Interaction and Retrieval, CHIIR 2017, Oslo, Norway, March 7–11, 2017, Ragnar Nordlie, Nils Pharo, Luanne Freund, Birger Larsen, and Dan Russel (Eds.). ACM, 117–126.
- Ram et al. (2023) Ori Ram, Yoav Levine, Itay Dalmedigos, Dor Muhlgay, Amnon Shashua, Kevin Leyton-Brown, and Yoav Shoham. 2023. In-Context Retrieval-Augmented Language Models. arXiv 2302.00083.
- Reimer et al. (2023) Jan Heinrich Reimer, Sebastian Schmidt, Maik Fröbe, Lukas Gienapp, Harrisen Scells, Benno Stein, Matthias Hagen, and Martin Potthast. 2023. The Archive Query Log: Mining Millions of Search Result Pages of Hundreds of Search Engines from 25 Years of Web Archives. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2023, Taipei, Taiwan, July 23–27, 2023, Hsin-Hsi Chen, Wei-Jou (Edward) Duh, Hen-Hsen Huang, Makoto P. Kato, Josiane Mothe, and Barbara Poblete (Eds.). ACM, 2848–2860.
- Renaud and Azzopardi (2012) Gareth Renaud and Leif Azzopardi. 2012. SCAMP: A Tool for Conducting Interactive Information Retrieval Experiments. In Information Interaction in Context: 2012, IIix’12, Nijmegen, The Netherlands, August 21–24, 2012, Jaap Kamps, Wessel Kraaij, and Norbert Fuhr (Eds.). ACM, 286–289.
- Reynolds and McDonell (2021) Laria Reynolds and Kyle McDonell. 2021. Prompt Programming for Large Language Models: Beyond the Few-Shot Paradigm. In CHI ’21: CHI Conference on Human Factors in Computing Systems, Virtual Event / Yokohama Japan, May 8–13, 2021, Extended Abstracts, Yoshifumi Kitamura, Aaron Quigley, Katherine Isbister, and Takeo Igarashi (Eds.). ACM, 314:1–314:7.
- Robinson and Wingate (2023) Joshua Robinson and David Wingate. 2023. Leveraging Large Language Models for Multiple Choice Question Answering. In The Eleventh International Conference on Learning Representations, ICLR 2023, Kigali, Rwanda, May 1–5, 2023. OpenReview.net, 28 pages.
- Roitero et al. (2018) Kevin Roitero, Eddy Maddalena, Gianluca Demartini, and Stefano Mizzaro. 2018. On Fine-Grained Relevance Scales. In The 41st International ACM SIGIR Conference on Research & Development in Information Retrieval, SIGIR 2018, Ann Arbor, MI, USA, July 08–12, 2018, Kevyn Collins-Thompson, Qiaozhu Mei, Brian D. Davison, Yiqun Liu, and Emine Yilmaz (Eds.). ACM, 675–684.
- Sakai (2023) Tetsuya Sakai. 2023. SWAN: A Generic Framework for Auditing Textual Conversational Systems. arXiv 2305.08290.
- Sakai et al. (2011) Tetsuya Sakai, Makoto P. Kato, and Young-In Song. 2011. Click the Search Button and be Happy: Evaluating Direct and Immediate Information Access. In Proceedings of the 20th ACM Conference on Information and Knowledge Management, CIKM 2011, Glasgow, United Kingdom, October 24–28, 2011, Craig Macdonald, Iadh Ounis, and Ian Ruthven (Eds.). ACM, 621–630.
- Sallam et al. (2023) Malik Sallam, Nesreen A Salim, B Ala’a, Muna Barakat, Diaa Fayyad, Souheil Hallit, Harapan Harapan, Rabih Hallit, Azmi Mahafzah, and B Ala’a. 2023. ChatGPT Output Regarding Compulsory Vaccination and COVID-19 Vaccine Conspiracy: A Descriptive Study at the Outset of a Paradigm Shift in Online Search for Information. Cureus 15, 2 (2023), 16 pages.
- Salton (1969) Gerard Salton. 1969. Interactive Information Retrieval. Technical Report. Cornell University.
- Sameki et al. (2016) Mehrnoosh Sameki, Aditya Barua, and Praveen K. Paritosh. 2016. Rigorously Collecting Commonsense Judgments for Complex Question-Answer Content. In Proceedings of the Third AAAI Conference on Human Computation and Crowdsourcing, HCOMP 2015, November 8–11, 2015, San Diego, California, USA, Volume 3, Elizabeth Gerber and Panos Ipeirotis (Eds.). AAAI Press, 26–33.
- Sander and Dietz (2021) David P. Sander and Laura Dietz. 2021. EXAM: How to Evaluate Retrieve-and-Generate Systems for Users Who Do Not (Yet) Know What They Want. In Proceedings of the Second International Conference on Design of Experimental Search & Information REtrieval Systems, Padova, Italy, September 15–18, 2021 (CEUR Workshop Proceedings, Vol. 2950), Omar Alonso, Stefano Marchesin, Marc Najork, and Gianmaria Silvello (Eds.). CEUR-WS.org, 136–146.
- Sanderson (2010) Mark Sanderson. 2010. Test Collection Based Evaluation of Information Retrieval Systems. Found. Trends Inf. Retr. 4, 4 (2010), 247–375.
- Schuff et al. (2022) Hendrik Schuff, Alon Jacovi, Heike Adel, Yoav Goldberg, and Ngoc Thang Vu. 2022. Human Interpretation of Saliency-based Explanation Over Text. In FAccT ’22: 2022 ACM Conference on Fairness, Accountability, and Transparency, Seoul, Republic of Korea, June 21–24, 2022. ACM, 611–636.
- Semnani et al. (2023) Sina J. Semnani, Violet Z. Yao, Heidi C. Zhang, and Monica S. Lam. 2023. WikiChat: A Few-Shot LLM-Based Chatbot Grounded with Wikipedia. arXiv 2305.14292.
- Shah et al. (2021) Darsh J. Shah, Lili Yu, Tao Lei, and Regina Barzilay. 2021. Nutri-bullets Hybrid: Consensual Multi-document Summarization. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, NAACL-HLT 2021, Online, June 6–11, 2021, Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tür, Iz Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou (Eds.). Association for Computational Linguistics, 5213–5222.
- Shi et al. (2023) Weijia Shi, Sewon Min, Michihiro Yasunaga, Minjoon Seo, Rich James, Mike Lewis, Luke Zettlemoyer, and Wen-tau Yih. 2023. REPLUG: Retrieval-Augmented Black-Box Language Models. arXiv 2301.12652.
- Shin et al. (2020) Taylor Shin, Yasaman Razeghi, Robert L. Logan IV, Eric Wallace, and Sameer Singh. 2020. AutoPrompt: Eliciting Knowledge from Language Models with Automatically Generated Prompts. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing, EMNLP 2020, Online, November 16–20, 2020, Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, 4222–4235.
- Snover et al. (2006) Matthew G. Snover, Bonnie J. Dorr, Richard M. Schwartz, Linnea Micciulla, and John Makhoul. 2006. A Study of Translation Edit Rate with Targeted Human Annotation. In Proceedings of the 7th Conference of the Association for Machine Translation in the Americas: Technical Papers, AMTA 2006, Cambridge, Massachusetts, USA, August 8–12, 2006. Association for Machine Translation in the Americas, 223–231.
- Song and Cohn (2011) Xingyi Song and Trevor Cohn. 2011. Regression and Ranking based Optimisation for Sentence Level MT Evaluation. In Proceedings of the Sixth Workshop on Statistical Machine Translation, WMT@EMNLP 2011, Edinburgh, Scotland, UK, July 30–31, 2011, Chris Callison-Burch, Philipp Koehn, Christof Monz, and Omar Zaidan (Eds.). Association for Computational Linguistics, 123–129.
- Sorensen et al. (2022) Taylor Sorensen, Joshua Robinson, Christopher Michael Rytting, Alexander Glenn Shaw, Kyle Jeffrey Rogers, Alexia Pauline Delorey, Mahmoud Khalil, Nancy Fulda, and David Wingate. 2022. An Information-Theoretic Approach to Prompt Engineering Without Ground Truth Labels. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), ACL 2022, Dublin, Ireland, May 22–27, 2022, Smaranda Muresan, Preslav Nakov, and Aline Villavicencio (Eds.). Association for Computational Linguistics, 819–862.
- Sun et al. (2023) Weiwei Sun, Lingyong Yan, Xinyu Ma, Shuaiqiang Wang, Pengjie Ren, Zhumin Chen, Dawei Yin, and Zhaochun Ren. 2023. Is ChatGPT Good at Search? Investigating Large Language Models as Re-Ranking Agents. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, EMNLP 2023, Singapore, December 6-10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, 14918–14937.
- Tay et al. (2022) Yi Tay, Vinh Tran, Mostafa Dehghani, Jianmo Ni, Dara Bahri, Harsh Mehta, Zhen Qin, Kai Hui, Zhe Zhao, Jai Prakash Gupta, Tal Schuster, William W. Cohen, and Donald Metzler. 2022. Transformer Memory as a Differentiable Search Index. In Advances in Neural Information Processing Systems 35: Annual Conference on Neural Information Processing Systems 2022, NeurIPS 2022, New Orleans, LA, USA, November 28 – December 9, 2022, Sanmi Koyejo, S. Mohamed, A. Agarwal, Danielle Belgrave, K. Cho, and A. Oh (Eds.). 13 pages.
- Thorne (2022) James Thorne. 2022. Data-Efficient Autoregressive Document Retrieval for Fact Verification. arXiv 2211.09388.
- Vakkari (2016) Pertti Vakkari. 2016. Searching as Learning: A Systematization based on Literature. J. Inf. Sci. 42, 1 (2016), 7–18.
- Wadden et al. (2020) David Wadden, Shanchuan Lin, Kyle Lo, Lucy Lu Wang, Madeleine van Zuylen, Arman Cohan, and Hannaneh Hajishirzi. 2020. Fact or Fiction: Verifying Scientific Claims. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing, EMNLP 2020, Online, November 16–20, 2020, Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, 7534–7550.
- Wang et al. (2022) Yujing Wang, Yingyan Hou, Haonan Wang, Ziming Miao, Shibin Wu, Qi Chen, Yuqing Xia, Chengmin Chi, Guoshuai Zhao, Zheng Liu, Xing Xie, Hao Sun, Weiwei Deng, Qi Zhang, and Mao Yang. 2022. A Neural Corpus Indexer for Document Retrieval. In Advances in Neural Information Processing Systems 35: Annual Conference on Neural Information Processing Systems 2022, NeurIPS 2022, New Orleans, LA, USA, November 28 – December 9, 2022, Sanmi Koyejo, S. Mohamed, A. Agarwal, Danielle Belgrave, K. Cho, and A. Oh (Eds.). 15 pages.
- White et al. (2023) Jules White, Quchen Fu, Sam Hays, Michael Sandborn, Carlos Olea, Henry Gilbert, Ashraf Elnashar, Jesse Spencer-Smith, and Douglas C. Schmidt. 2023. A Prompt Pattern Catalog to Enhance Prompt Engineering with ChatGPT. arXiv 2302.11382.
- Wikimedia Foundation (2023) Wikimedia Foundation. 2023. Wikipedia: Verifiability, not Truth. https://web.archive.org/web/20230627143645/https://en.wikipedia.org/wiki/Wikipedia:Verifiability,_not_truth. Accessed: 2023-06-27.
- Wilson (2011) Max L. Wilson. 2011. Interfaces for Information Retrieval. In Interactive Information Seeking, Behaviour and Retrieval, Ian Ruthven and Diane Kelly (Eds.). Facet Publishing, 139–170.
- Wu et al. (2023) Zhijing Wu, Jiaxin Mao, Kedi Xu, Dandan Song, and Heyan Huang. 2023. A Passage-Level Reading Behavior Model for Mobile Search. In Proceedings of the ACM Web Conference 2023, WWW 2023, Austin, TX, USA, 30 April 2023 – 4 May 2023, Ying Ding, Jie Tang, Juan F. Sequeda, Lora Aroyo, Carlos Castillo, and Geert-Jan Houben (Eds.). ACM, 3236–3246.
- Yang et al. (2023) Chengrun Yang, Xuezhi Wang, Yifeng Lu, Hanxiao Liu, Quoc V. Le, Denny Zhou, and Xinyun Chen. 2023. Large Language Models as Optimizers. arXiv 2309.03409.
- Yang (2017) Ziying Yang. 2017. Relevance Judgments: Preferences, Scores and Ties. In Proceedings of the 40th International ACM SIGIR Conference on Research and Development in Information Retrieval, Shinjuku, Tokyo, Japan, August 7–11, 2017, Noriko Kando, Tetsuya Sakai, Hideo Joho, Hang Li, Arjen P. de Vries, and Ryen W. White (Eds.). ACM, 1373.
- Yu et al. (2023) Wenhao Yu, Dan Iter, Shuohang Wang, Yichong Xu, Mingxuan Ju, Soumya Sanyal, Chenguang Zhu, Michael Zeng, and Meng Jiang. 2023. Generate Rather Than Retrieve: Large Language Models are Strong Context Generators. In The Eleventh International Conference on Learning Representations, ICLR 2023, Kigali, Rwanda, May 1–5, 2023. OpenReview.net, 27 pages.
- Yuan et al. (2021) Weizhe Yuan, Graham Neubig, and Pengfei Liu. 2021. BARTScore: Evaluating Generated Text as Text Generation. In Advances in Neural Information Processing Systems 34: Annual Conference on Neural Information Processing Systems 2021, NeurIPS 2021, December 6–14, 2021, virtual, Marc’Aurelio Ranzato, Alina Beygelzimer, Yann N. Dauphin, Percy Liang, and Jennifer Wortman Vaughan (Eds.). 27263–27277.
- Yue et al. (2023) Xiang Yue, Boshi Wang, Ziru Chen, Kai Zhang, Yu Su, and Huan Sun. 2023. Automatic Evaluation of Attribution by Large Language Models. In Findings of the Association for Computational Linguistics: EMNLP 2023, Singapore, December 6–10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, 4615–4635.
- Zamani et al. (2020) Hamed Zamani, Bhaskar Mitra, Everest Chen, Gord Lueck, Fernando Diaz, Paul N. Bennett, Nick Craswell, and Susan T. Dumais. 2020. Analyzing and Learning from User Interactions for Search Clarification. In Proceedings of the 43rd International ACM SIGIR conference on research and development in Information Retrieval, SIGIR 2020, Virtual Event, China, July 25–30, 2020, Jimmy X. Huang, Yi Chang, Xueqi Cheng, Jaap Kamps, Vanessa Murdock, Ji-Rong Wen, and Yiqun Liu (Eds.). ACM, 1181–1190.
- Zhang and Pradeep (2023) Dake Zhang and Ronak Pradeep. 2023. ReadProbe: A Demo of Retrieval-Enhanced Large Language Models to Support Lateral Reading. arXiv 2306.07875.
- Zhang et al. (2021a) Edwin Zhang, Sheng-Chieh Lin, Jheng-Hong Yang, Ronak Pradeep, Rodrigo Frassetto Nogueira, and Jimmy Lin. 2021a. Chatty Goose: A Python Framework for Conversational Search. In SIGIR ’21: The 44th International ACM SIGIR Conference on Research and Development in Information Retrieval, Virtual Event, Canada, July 11–15, 2021, Fernando Diaz, Chirag Shah, Torsten Suel, Pablo Castells, Rosie Jones, and Tetsuya Sakai (Eds.). ACM, 2521–2525.
- Zhang et al. (2020) Tianyi Zhang, Varsha Kishore, Felix Wu, Kilian Q. Weinberger, and Yoav Artzi. 2020. BERTScore: Evaluating Text Generation with BERT. In 8th International Conference on Learning Representations, ICLR 2020, Addis Ababa, Ethiopia, April 26–30, 2020. OpenReview.net, 43 pages.
- Zhang et al. (2021b) Zijian Zhang, Koustav Rudra, and Avishek Anand. 2021b. Explain and Predict, and then Predict Again. In WSDM ’21, The Fourteenth ACM International Conference on Web Search and Data Mining, Virtual Event, Israel, March 8–12, 2021, Liane Lewin-Eytan, David Carmel, Elad Yom-Tov, Eugene Agichtein, and Evgeniy Gabrilovich (Eds.). ACM, 418–426.
- Zhao et al. (2023) Ruochen Zhao, Xingxuan Li, Yew Ken Chia, Bosheng Ding, and Lidong Bing. 2023. Can ChatGPT-like Generative Models Guarantee Factual Accuracy? On the Mistakes of New Generation Search Engines. arXiv 2304.11076.
- Zhao et al. (2019) Wei Zhao, Maxime Peyrard, Fei Liu, Yang Gao, Christian M. Meyer, and Steffen Eger. 2019. MoverScore: Text Generation Evaluating with Contextualized Embeddings and Earth Mover Distance. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing, EMNLP-IJCNLP 2019, Hong Kong, China, November 3–7, 2019, Kentaro Inui, Jing Jiang, Vincent Ng, and Xiaojun Wan (Eds.). Association for Computational Linguistics, 563–578.
- Zheng et al. (2012) Wei Zheng, Xuanhui Wang, Hui Fang, and Hong Cheng. 2012. Coverage-Based Search Result Diversification. Information Retrieval 15, 5 (2012), 433–457.
- Zheng et al. (2019) Yukun Zheng, Jiaxin Mao, Yiqun Liu, Zixin Ye, Min Zhang, and Shaoping Ma. 2019. Human Behavior Inspired Machine Reading Comprehension. In Proceedings of the 42nd International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2019, Paris, France, July 21–25, 2019, Benjamin Piwowarski, Max Chevalier, Éric Gaussier, Yoelle Maarek, Jian-Yun Nie, and Falk Scholer (Eds.). ACM, 425–434.
- Zhou et al. (2023) Yujia Zhou, Jing Yao, Zhicheng Dou, Ledell Wu, and Ji-Rong Wen. 2023. DynamicRetriever: A Pre-trained Model-based IR System Without an Explicit Index. Mach. Intell. Res. 20, 2 (2023), 276–288.
- Zhu et al. (2009) Zhemin Zhu, Delphine Bernhard, and Iryna Gurevych. 2009. A Multi-Dimensional Model for Assessing the Quality of Answers in Social Q&A Sites. In Proceedings of the 14th International Conference on Information Quality, ICIQ 2009, Hasso Plattner Institute, University of Potsdam, Germany, November 7–8 2009, Paul L. Bowen, Ahmed K. Elmagarmid, Hubert Österle, and Kai-Uwe Sattler (Eds.). HPI/MIT, 264–265.
- Zhuang et al. (2023) Honglei Zhuang, Zhen Qin, Kai Hui, Junru Wu, Le Yan, Xuanhui Wang, and Michael Berdersky. 2023. Beyond Yes and No: Improving Zero-Shot LLM Rankers via Scoring Fine-Grained Relevance Labels. arXiv 2310.14122.
- Zhuang et al. (2022) Shengyao Zhuang, Houxing Ren, Linjun Shou, Jian Pei, Ming Gong, Guido Zuccon, and Daxin Jiang. 2022. Bridging the Gap Between Indexing and Retrieval for Differentiable Search Index with Query Generation. arXiv 2206.10128.
- Zhuang and Zuccon (2021) Shengyao Zhuang and Guido Zuccon. 2021. Fast Passage Re-ranking with Contextualized Exact Term Matching and Efficient Passage Expansion. arXiv 2108.08513.
- Ziems et al. (2023) Noah Ziems, Wenhao Yu, Zhihan Zhang, and Meng Jiang. 2023. Large Language Models are Built-in Autoregressive Search Engines. In Findings of the Association for Computational Linguistics: ACL 2023, Toronto, Canada, July 9–14, 2023, Anna Rogers, Jordan L. Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, 2666–2678.